### v2

# disposition

Abstract
It is well-known that scalarization techniques (e.g., in the sense of Gerstewitz; Kasimbeyli; Pascoletti and Serafini; Zaffaroni) are useful for generating (weakly, properly) efficient solutions of vector optimization problems. One recognized approach is the conic scalarization method in vector optimization in real normed spaces proposed by Kasimbeyli (2010, SIAM J Optim 20), which is based on augmented dual cones and Bishop-Phelps type (norm-linear) scalarizing functions. In this paper, we present new results on cone separation in real topological-linear spaces by using Bishop-Phelps type separating cones / separating seminorm-linear functions. Moreover, we study some extensions of known scalarization results in vector optimization (in the sense of Eichfelder; Gerstewitz; Jahn; Kasimbeyli; Pascoletti and Serafini). On this basis, we propose a Bishop-Phelps type scalarization method for vector optimization problems in real topological-linear spaces, which can be seen as an extension of Kasimbeyli’s conic scalarization method in real normed spaces. Within this framework, we derive new Bishop-Phelps type scalarization results for the concepts of weak efficiency and different types of proper efficiency.
Abstract
Keywords: Vector optimization; (proper, weak) efficiency; Bishop-Phelps cone; augmented dual cone; Bishop-Phelps type scalarization; conic scalarization
Mathematics subject classifications (MSC 2010): 90C48, 90C29, 06F20, 52A05
1 Introduction
The field of vector optimization in real topological-linear spaces is concerned with the minimization of a vector-valued function $f:X\to Y$ acting between real topological-linear spaces $X$ and $Y$ with respect to a convex cone $C\subseteq Y$ over a nonempty feasible set $\Omega\subseteq X$. A review of the existing literature reveals that many papers focus on the (finite dimensional) normed setting (i.e., $Y$ is a normed space), see e.g. [16], [8, 9], [12, 13], [17], [32], [45, 46], [53], [58], [57, Ch. 7 and Sec. 15.2], and references therein. In particular, the analysis includes real normed spaces $Y$ where the convex cone $C$ is of Bishop-Phelps type (named after Bishop and Phelps [2], see also [16], [15], [28, 29], [36, 37], [35],[41, 43], [45, 46]). The importance of the class of Bishop-Phelps cones can be emphasized by the fact that a nontrivial, closed, convex cone in a real normed space is representable as a Bishop-Phelps cone if and only if it is well-based (see Example 2.4 for details). We are primarily interested on the more general setting of a real (possibly infinite dimensional and not normed) topological-linear space $Y$ motivated by optimization problems under uncertainty and problems from mathematical finance. Especially, robust counterparts to optimization problems under uncertainty can be formulated as vector optimization problems with an infinite dimensional topological-linear image space (see Göpfert et al. [23, Ch. 6] and references therein). Furthermore, in arbitrage theory (see the class of financial market models in Marohn and Tammer [51, 52], and Riedel [56]), problems in the setting of linear spaces are considered. Some further works in vector optimization that focus on this setting are [5], [20, 21], [22], [23], [30], [34], [25], [42, 43], [57].
The main focus in our paper is on the computation of efficient / weakly efficient / properly efficient solutions of the given vector optimization problem via scalarization (based on Bishop-Phelps type cones), where the proper efficiency concepts considered in this paper cover certain known concepts of proper efficiency from the literature (e.g., in the sense of Benson [1]; Borwein [3, 4]; Henig [38]; Hurwicz [40]). It is well established that scalarization methods are useful for solving vector optimization problems. A notable feature of such methods is the ability to compute weakly efficient solutions by solving scalar optimization problems, which are often parameter-dependent, instead of the original vector optimization problem. Moreover, using scalarization, one can derive necessary and sufficient optimality conditions for the initial vector optimization problem, using generalized differentiation concepts for the new, scalarized objective function (see, e.g., [24], [32], [10], [33], [8], [34], [9]). The application of the scalarization methods for Henig dilating cones (enlarging convex cones) of the convex cone $C$ yields solutions that are properly efficient in the sense of Henig [38]. The relationships of the Henig proper efficiency concept to other known proper efficiency concepts have been well studied in the literature (see, e.g., [5, Sec. 2.4.3], [25, Ch. 3], [26], [49, Sec. 2.4] and references therein).
We will now recall some well-known scalarization methods in vector optimization. Of course, the linear scalarization method in vector optimization (respectively, the weighted sum method in multiobjective optimization) is the best known technique for generating weakly / properly efficient solutions of vector (multiobjective) optimization problems. It is known that this method only works well under cone-convexity type assumptions (see, e.g., [27]). For nonconvex vector optimization problems, there are some other scalarization methods frequently used in the literature that we recall below. In the framework of real (possibly infinite dimensional) normed spaces, the conic scalarization method by Kasimbeyli [45, 46] (see also [16]), the Drummond–Svaiter scalarization method [7] (see also [6]) and the Zaffaroni scalarization method [58] (see also [14, Sec. 5.1]) based on the Hiriart-Urruty function [39] (also known as oriented/signed distance function) are known to be useful, where the scalarizing function of the first two methods involves the underlying norm of the space $Y$. Considering the more general framework of real topological-linear spaces, only a few methods (besides the linear scalarization method) are known to be valid, such as the Gerstewitz scalarization method [20, 21] (see also [22], [49], [57], [23]) and the Pascoletti–Serafini scalarization method [53] (see also [12, 13], [57, Sec. 6.2.2]). Certain well-known scalarization methods in vector and multiobjective optimization (e.g. linear scalarization method / weighted sum method, $\varepsilon$-constraint method, hybrid method, Benson method, weighted Chebyshev norm method, and other methods; see [57, Sec. 7.3]) are special cases of these general scalarization methods. Some further comparison studies of scalarization methods in vector optimization are given in [6], [48], [12, Ch. 2], while generalizations of scalarization methods to vector and set optimization involving variable domination structures are also studied in the recent literature (see, e.g., [35], [15], [14], [43], [57, Ch. 10 and 11] and references therein).
As mentioned above, one recognized approach in vector optimization in a real normed space $(Y,||\cdot||)$ is the conic scalarization method in vector optimization proposed by Kasimbeyli [45], which is based on the following parameterized scalar optimization problem:
$\begin{cases}x^{*}(f(x)-a)+\alpha||f(x)-a||\to\min\\ x\in\Omega\end{cases}$ (1.1)
with a pair $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$ (respectively, $(x^{*},\alpha)$ belongs to the so-called augmented dual cone of $C$; see Section 2.3 for details) and a given reference point $a\in Y$. Several relationships between (weakly, properly) efficient solutions of the original vector problem and solutions of scalar problems of type (1.1) are known (see [16], [17], [31], [45, 46]). In this paper, we propose a Bishop-Phelps type scalarization method for vector optimization problems in real topological-linear spaces, which can be seen as an extension of Kasimbeyli’s scalarization method in real normed spaces. By taking a look on (1.1) one can see that Kasimbeyli’s method is based on Bishop-Phelps type norm-linear scalarizing functions (i.e., functions that are a nonnegative weighted sum of the norm |||||$\cdot$||| | ⋅ | | with a linear continuous functional $x^{*}$, see also Zaffaroni [59]), where the involving norm |||||$\cdot$||| | ⋅ | | is exactly the underlying norm in the real normed space $Y$. In our Bishop-Phelps type scalarization method, we consider Bishop-Phelps type seminorm-linear scalarizing functions (i.e., functions that are a nonnegative weighted sum of a seminorm $\psi$ with a linear continuous functional $x^{*}$), where the seminorm involved (namely $\psi$) must not have anything to do with the underlying topology of the real (locally convex) topological-linear space (e.g., the seminorm must not be part of a family of seminorms that generates this topology).
The rest of the paper is organized as follows:
In Section 2, after presenting some basics in real topological-linear spaces including properties of seminorms, we study properties of cones and their bases, focusing in particular on augmented dual cones, Bishop-Phelps type cones and zero level sets of Bishop-Phelps type (seminorm-linear) functions. We then introduce the solution concepts of vector optimization necessary for our work.
Section 3 is devoted to the study of scalarization methods in vector optimization which are based on cone monotone and cone representing scalarizing functions. Beside classical scalarization results (in the sense of Eichfelder [12, 13]; Gerstewitz [20, 21]; Jahn [43]; Pascoletti and Serafini [53]), we establish some new scalarization results (namely Propositions 3.11, 3.14 and 3.17) for vector optimization problems in real topological-linear spaces, which extend the classical ones.
In Section 4, we present the key tool to prove our main Bishop-Phelps type scalarization results in vector optimization, namely the technique of nonlinear separation of two (not necessarily convex) cones by using Bishop-Phelps type separating cones / separating (seminorm-linear) functions. Besides presenting some classical nonlinear cone separation results, we also reflect on the separation conditions involved in both real topological-linear spaces and real normed spaces. Propositions 4.1 and 4.11 give some new insights into the relationships between the classical separation condition (the intersection of two cones is the singleton set $\{0\}$) and the separation conditions used in the nonlinear separation approach based on Bishop-Phelps type cones / functions discussed in [16], [28, 29], [45, 46]. We are able to deduce some new variants of nonlinear cone separation theorems involving the classical separation condition (see Corollaries 4.16 and 4.17).
The main Section 5 is dedicated to Bishop-Phelps type scalarization for vector optimization in real topological-linear spaces. The parameterized scalarization problems (in the sense of Kasimbeyli [45]; Gerstewitz [20, 21]; Jahn [43]; Pascoletti and Serafini [53]) are formulated using Bishop-Phelps type cone representing (seminorm-linear) functions. In addition to the basic scalarization results for the original vector optimization problem (roughly speaking, solutions of the scalar problems are solutions of the vector problem), we also point out (for the concept of efficiency in Proposition 5.12 and for the concept of weak efficiency in Proposition 5.14) the relationships between the original vector problem (based on the convex cone $C$), an updated vector problem based on an enlarging / dilating cone of Bishop-Phelps type, and scalar problems involving (Bishop-Phelps type) cone representing and cone monotone scalarizing functions (e.g., seminorm-linear scalarizing functions or Gerstewitz [20, 21] scalarizing functions). Then, using nonlinear cone separation results (from Section 4) and Bishop-Phelps type scalarization results (Propositions 5.12 and 5.14), we are able to present new scalarization theorems (Theorem 5.19 for weak efficiency, Theorems 5.25, 5.38 and 5.40 for proper efficiency) and their corollaries, which show that any (weakly, properly) efficient solution of the original vector problem is a solution of the above mentioned (parameterized) scalar problems for certain choices of parameters.
Finally, in Section 6 we present a conclusion and an outlook for further research.
2 Preliminaries
2.1 Basics in topological-linear spaces
Throughout the paper, we assume that $Y$ is a real topological-linear space. As usual, $Y^{*}$ denotes the topological dual space of $Y$. In certain situations we will also consider a more specific setting where $(Y,||\cdot||)$ is a real normed space. Consistently, the set of nonnegative real numbers is denoted by $\mathbb{R}_{+}$, while $\mathbb{P}:=\mathbb{R}_{++}$ denotes the set of positive real numbers. Given a set $\Omega\subseteq Y$, we denote by ${\mathrm{c}onv}\,\Omega$ the convex hull of $\Omega$. Moreover, we denote by ${\mathrm{c}l}\,\Omega$, ${\mathrm{b}d}\,\Omega$, and ${\mathrm{i}nt}\,\Omega$ the closure, the boundary, and the interior of $\Omega$ with respect to the underlying topology in $Y$. The algebraic interior (the core) of $\Omega$ is defined by $\operatorname{cor}\,\Omega:=\{x\in\Omega\mid\forall\,v\in Y\;\exists\,\varepsilon>0:\;x+[0,\varepsilon]\cdot v\subseteq\Omega\}.$ It is known that $\operatorname{int}\,\Omega\subseteq\operatorname{cor}\,\Omega\subseteq\Omega\subseteq\operatorname{cl}\,\Omega=(\operatorname{int}\,\Omega)\cup\operatorname{bd}\,\Omega,$ and if $\operatorname{int}\,\Omega\neq\emptyset$ (i.e., $\Omega$ is solid) and convex, then $\operatorname{cor}\,\Omega=\operatorname{int}\,\Omega$.
Given a seminorm $\psi:Y\to\operatorname{\mathbb{R}}$, we define the $\psi$-unit ball by $\operatorname{\mathbb{B}}_{\psi}:=\{y\in Y\mid\psi(y)\leq 1\}$ and the $\psi$-unit sphere by $\operatorname{\mathbb{S}}_{\psi}:=\{y\in Y\mid\psi(y)=1\}.$
Remark 2.1
Suppose that $Y$ is a real topological-linear space and $\psi:Y\to\mathbb{R}$ is a seminorm. The following assertions hold:
  • $1^{\circ}$.
    $\psi$ is upper semicontinuous $\iff$ the set $\{y\in Y\mid\psi(y)<1\}$ is open $\iff$ $\psi$ is continuous at the origin $\iff$ $\psi$ is continuous.
  • $2^{\circ}$.
    $\psi$ is lower semicontinuous $\iff$ $\operatorname{\mathbb{B}}_{\psi}$ is closed.
Remark that ${\mathrm{k}er}\,\psi=\{0\}$ if and only if $\psi$ is a norm, and the singleton set $\{0\}$ is closed if and only if the topological-linear space $Y$ is Hausdorff. Moreover, it is known (see, e.g., Gelfand [18, 19], Eberlein [11]) that if $Y$ is a Banach space and $\psi:Y\to\mathbb{R}$ is a lower semicontinuous seminorm (alternatively, the seminorm $\psi$ has the Baire property), then $\psi$ is continuous.
2.2 Cones and their bases
A set $K\subseteq Y$ is called a cone if $0\in K=\operatorname{\mathbb{R}}_{+}\cdot K$. A cone $K\subseteq Y$ is said to be convex if $K+K=K$ (or equivalently, $K$ is a convex set); proper or nontrivial if $\{0\}\neq K\neq Y$; pointed if $\ell(K):=K\cap(-K)=\{0\}$. Given a convex cone $K\subseteq Y$ in the real topological-linear space $Y$, the set $\ell(K)$ is known as the lineality space of $K$. In this case, $K$ is a linear subspace of $Y$ if and only if $K=\ell(K)$, and moreover, $\operatorname{int}\,K=K+\operatorname{int}\,K$ and $K\setminus\ell(K)=K+(K\setminus\ell(K))$. If $K\subseteq Y$ is a cone, then $K\neq Y\iff 0\notin\operatorname{cor}\,K\iff 0\notin\operatorname{int}\,K,$ and if further $K$ is convex, then (see also [27, Lem. 2.10]) $K\neq Y\iff\ell(K)\cap\operatorname{int}\,K=\emptyset.$
The topological dual cone of a cone $K\subseteq Y$ is given by $K^{+}:=\{y^{*}\in Y^{*}\mid\forall\,k\in K:\;y^{*}(k)\geq 0\}.$ Furthermore, the subset $K^{\#}:=\{y^{*}\in Y^{*}\mid\forall\,k\in K\setminus\{0\}:\;y^{*}(k)>0\}$ of $K^{+}$ is of special interest. It is important to mention that both sets $K^{+}$ and $K^{\#}$ are convex for any (not necessarily convex) cone $K\subseteq Y$. Moreover, one has
$K^{+}=(\operatorname{conv}\,K)^{+}=(\operatorname{cl}(\operatorname{conv}\,K))^{+}\text{ and }(\operatorname{cl}(\operatorname{conv}\,K))^{\#}\subseteq K^{\#}=(\operatorname{conv}\,K)^{\#},$ (2.1)
but the inclusion $(\operatorname{cl}(\operatorname{conv}\,K))^{\#}\subseteq K^{\#}$ can be strict (see Göpfert et al. [23, p. 55]).
In the following definition, we recall a general base concept for cones in real topological-linear spaces (cf. Göpfert et al. [23, Def. 2.1.42]).
Definition 2.2
Consider a nontrivial cone $K\subseteq Y$. A set $B\subseteq K$ is called a base for $K$, if $B$ is a nonempty set, and $K=\operatorname{\mathbb{R}}_{+}\cdot B$ with $0\notin\operatorname{cl}\,B$. Moreover, $K$ is said to be well-based if there exists a bounded, convex base of $K$.
Remark 2.3
Taking into account (2.1), for any nontrivial cone $K\subseteq Y$ in a real normed space $Y$, it is well-known (see [23, Prop. 2.2.23 and 2.2.32]) that
$\operatorname{conv}\,K\text{ has a convex base}\iff K^{\#}\neq\emptyset,$ (2.2)
and
$\operatorname{conv}\,K\text{ is well-based}\iff\operatorname{cl}(\operatorname{conv}\,K)\text{ is well-based}\iff\operatorname{int}\,K^{+}\neq\emptyset.$ (2.3)
Moreover, if $Y$ has finite dimension and $K$ is closed, then $K^{\#}=\operatorname{int}\,K^{+}$ (see [29, Th. 2.1(4), Rem. 2.6]), hence all the conditions involved in (2.2) and (2.3) are equivalent.
Example 2.4
Assume that $Y$ is a real normed space with the underlying norm $||\cdot||:Y\to\mathbb{R}$. For any $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$, a Bishop-Phelps cone (in the sense of Bishop and Phelps [2]) is defined by $C_{||\cdot||}(x^{*},\alpha):=\{y\in Y\mid x^{*}(y)\geq\alpha||y||\}$. It is known that Bishop-Phelps cones have a lot of useful properties and there are interesting applications in variational analysis and optimization (see, e.g., Ha and Jahn [36, 37], Phelps [55]). Any Bishop-Phelps cone $C_{||\cdot||}(x^{*},\alpha)$ is a closed, pointed, convex cone. If $||x^{*}||_{*}>\alpha$ (where $||\cdot||_{*}:Y^{*}\to\mathbb{R}$ denotes the dual norm of |||||$\cdot$||| | ⋅ | |), then $C_{||\cdot||}(x^{*},\alpha)$ is nontrivial and $C_{||\cdot||}^{>}(x^{*},\alpha):=\{y\in Y\mid x^{*}(y)>\alpha||y||\}=\operatorname{cor}\,C_{||\cdot||}(x^{*},\alpha)=\operatorname{int}\,C_{||\cdot||}(x^{*},\alpha)\neq\emptyset$. Note that for $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ with $||x^{*}||_{*}<\alpha$ we have $C_{||\cdot||}(x^{*},\alpha)=\{0\}$, while for $||x^{*}||_{*}\leq\alpha$ we have $C_{||\cdot||}^{>}(x^{*},\alpha)=\emptyset$. It is worth noting that every nontrivial, closed, convex cone $K\subseteq Y$ in a normed space $Y$ is representable as a Bishop-Phelps cone (i.e., there is a pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ and an equivalent norm $||\cdot||^{\prime}\sim||\cdot||$ such that $K=C_{||\cdot||^{\prime}}(x^{*},\alpha)$) if and only if $K$ is well-based (see the proof of Proposition 4.11). Note that any nontrivial, closed, pointed, convex cone $K$ in a finite dimensional normed space $Y$ is well-based, since in this setting $K$ is well-based if and only if $K^{\#}\neq\emptyset$ (taking into account Remark 2.3), and the latter condition $K^{\#}\neq\emptyset$ is true by the well-known Krein-Rutman theorem (see, e.g., [49, Th. 2.4.7]).
Moreover, we recall another (algebraic) base concept for cones in linear spaces (cf. Göpfert et al. [23, Def. 2.1.14], Jahn [42, Def. 1.10 (d)]).
Definition 2.5
Consider a nontrivial cone $K\subseteq Y$. A set $B\subseteq K\setminus\{0\}$ is called an algebraic base for $K$, if $B$ is a nonempty set, and every $x\in K\setminus\{0\}$ has a unique representation of the form $x=\lambda b$ for some $\lambda>0$ and some $b\in B$.
Remark 2.6
If $B$ is base in the sense of Definition 2.2 or Definition 2.5 for the nontrivial cone $K\subseteq Y$, then $K=\operatorname{\mathbb{R}}_{+}\cdot B$ and $K\setminus\{0\}=\operatorname{\mathbb{P}}\cdot B$.
Our focus will be on a specific type of base in real topological-linear spaces, which we will recall in the following definition.
Definition 2.7
Consider a seminorm $\psi:Y\to\operatorname{\mathbb{R}}$ and a cone $K\subseteq Y$. Define
$B_{K}:=B_{K}(\psi):=\{x\in K\mid\psi(x)=1\}=K\cap\operatorname{\mathbb{S}}_{\psi}$.
If $K$ is nontrivial and $\psi$ is positive on $K\setminus\{0\}$, then $B_{K}$ is called a normlike-base for $K$. If $\psi$ is a norm $||\cdot||:Y\to\operatorname{\mathbb{R}}$, then $B_{K}(||\cdot||)$ is called a norm-base for $K$. Note that $B_{K}(\psi)$ is a base in the sense of Definition 2.5, and if $\psi$ is continuous and $K$ is closed, then it is a base also in the sense of Definition 2.2.
2.3 Augmented dual cones and Bishop-Phelps type cones
Assume that $Y$ is a topological-linear space, $K\subseteq Y$ is a nontrivial cone, and $\psi:Y\to\mathbb{R}$ is a seminorm. For any $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$, we consider the Bishop-Phelps type (seminorm-linear) function $\varphi_{x^{*},\alpha}:Y\to\mathbb{R}$ defined by
$\varphi_{x^{*},\alpha}(y):=x^{*}(y)+\alpha\psi(y)\quad\text{for all }y\in Y,$
as used in [28]. Note that Bishop-Phelps type (norm-linear) functions in a normed setting were studied for instance by García-Castaño, Melguizo-Padial and Parzanese [16], Gasimov (Kasimbeyli) [17], Günther, Khazayel and Tammer [29], Guo and Hang [31], Ha [35], Jahn [43], Kasimbeyli [45, 46], and Zaffaroni [59].
Following the definitions of so-called augmented dual cones by Gasimov [17], Kasimbeyli [45] (see also [28] for some generalizations), we consider a (generalized) augmented dual cone by
$K^{a+}(\psi):=\{(x^{*},\alpha)\in K^{+}\times\mathbb{R}_{+}\mid\forall\,y\in K:\;x^{*}(y)-\alpha\psi(y)\geq 0\}$
as well the following sets
$K^{a\circ}(\psi)$ :=$:=\{(x^{*},\alpha)\in(K^{+}\setminus\{0\})\times\mathbb{R}_{+}\mid\forall\,y\in\operatorname{int}\,K:\;x^{*}(y)-\alpha\psi(y)>0\},$
$K^{a\#}(\psi)$ :=$:=\{(x^{*},\alpha)\in K^{\#}\times\mathbb{R}_{+}\mid\forall\,y\in K\setminus\{0\}:\;x^{*}(y)-\alpha\psi(y)>0\}.$
If no danger of confusion arises, we will use the notations $K^{a+}$, $K^{a\circ}$ and $K^{a\#}$.
Remark 2.8
Note that $(\operatorname{conv}\,K)^{a+}=K^{a+}$ and $(\operatorname{conv}\,K)^{a\#}=K^{a\#}$, and if $\psi$ is lower semicontinuous, then $(\operatorname{cl}(\operatorname{conv}\,K))^{a+}=K^{a+}$. It is a simple observation that for a normlike-base $B_{K}=B_{K}(\psi)$ of $K$ we have
$K^{a+}$ =$=\{(x^{*},\alpha)\in K^{+}\times\mathbb{R}_{+}\mid\forall\,y\in B_{K}:\;x^{*}(y)\geq\alpha\},$
$K^{a\#}$ =$=\{(x^{*},\alpha)\in K^{\#}\times\mathbb{R}_{+}\mid\forall\,y\in B_{K}:\;x^{*}(y)>\alpha\}.$
Furthermore, it is easy to observe that $K^{a\#}\subseteq K^{a\circ}\cup K^{a+}$, and if $K$ is solid and $\psi$ is lower semicontinuous, then $K^{a\circ}\subseteq K^{a+}$.
In the following definition, we introduce two classes of functions that satisfy cone (interior) representation properties (cf. Ha [35, Sec. 3.1], Jahn [43, Th. 3.1]).
Definition 2.9
The functional $\varphi:Y\to\mathbb{R}$ is
  • $-K$ representing if $-K=\{y\in Y\mid\varphi(y)\leq 0\}.$
  • strictly $-K$ representing if $\varphi$ is $-K$ representing and $-\operatorname{int}\,K=\{y\in Y\mid\varphi(y)<0\}.$
Remark 2.10
It is known that several scalarizing functions in vector optimization (e.g., in the sense of Gerstewitz [20, 21]; Hiriart-Urruty [39] / Zaffaroni [58, Prop. 3.2]) admit cone representation properties. Moreover, for certain topological-linear spaces $Y$, which are preordered by a (solid) convex cone $K\subseteq Y$, one can state $-K$ representing (strictly $-K$ representing) functionals (see, e.g., Günther et al. [30, Sec. 4.1], Ha [35], and Jahn [43]).
For any $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$, the sublevel set of the Bishop-Phelps type function $\varphi_{x^{*},\alpha}$ at level zero is defined by
$S^{\leq}_{\psi}(x^{*},\alpha)$ :=$:=\{y\in Y\mid\varphi_{x^{*},\alpha}(y)=x^{*}(y)+\alpha\psi(y)\leq 0\},$
while a Bishop-Phelps type cone is defined by
$C_{\psi}(x^{*},\alpha):=\{y\in Y\mid x^{*}(y)\geq\alpha\psi(y)\}.$
Remark that, in case $\psi=||\cdot||$ and $\alpha>0$, this reduces to the usual Bishop-Phelps cone (see Example 2.4). Moreover, the sets
$S^{<}_{\psi}(x^{*},\alpha)$ :=$:=\{y\in Y\mid\varphi_{x^{*},\alpha}(y)=x^{*}(y)+\alpha\psi(y)<0\},$
$C^{>}_{\psi}(x^{*},\alpha)$ :=$:=\{y\in Y\mid x^{*}(y)>\alpha\psi(y)\}$
are of interest. In our work we will focus in particular on Bishop-Phelps type functions, since they are (strictly) $-K$ representing functions if $K$ is a Bishop-Phelps type cone. Let us first note that the sets $K^{a+}$, $K^{a\circ}$ and $K^{a\#}$ can be written as
$K^{a+}$ =$=\{(x^{*},\alpha)\in K^{+}\times\mathbb{R}_{+}\mid K\subseteq C_{\psi}(x^{*},\alpha)\},$
$K^{a\circ}$ =$=\{(x^{*},\alpha)\in(K^{+}\setminus\{0\})\times\mathbb{R}_{+}\mid\operatorname{int}\,K\subseteq C^{>}_{\psi}(x^{*},\alpha)\},$
$K^{a\#}$ =$=\{(x^{*},\alpha)\in K^{\#}\times\mathbb{R}_{+}\mid K\setminus\{0\}\subseteq C^{>}_{\psi}(x^{*},\alpha)\}.$
In the following lemmata, we collect some important properties of Bishop-Phelps type cones and functions. For the first lemma, the proof is straightforward and we omit it.
Lemma 2.11
Consider $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$. The following assertions hold:
  • $1^{\circ}$.
    Cone property: $0\in C_{\psi}(x^{*},\alpha),\operatorname{\mathbb{P}}\cdot C_{\psi}(x^{*},\alpha)=C_{\psi}(x^{*},\alpha)$ and $\operatorname{\mathbb{P}}\cdot C^{>}_{\psi}(x^{*},\alpha)=C^{>}_{\psi}(x^{*},\alpha)$.
  • $2^{\circ}$.
    Convexity: $C_{\psi}(x^{*},\alpha)$ and $C^{>}_{\psi}(x^{*},\alpha)$ are convex.
  • $3^{\circ}$.
    Closedness: If $\varphi_{x^{*},\alpha}$ (respectively, $\psi$) is lower semicontinuous, then $C_{\psi}(x^{*},\alpha)$ is closed.
  • $4^{\circ}$.
    Pointedness: $\ell(C_{\psi}(x^{*},\alpha))=\{y\in Y\mid x^{*}(y)=0,\alpha\psi(y)=0\}$. If $\psi$ is a norm and $\alpha>0$, then $C_{\psi}(x^{*},\alpha)$ is pointed. If $(x^{*},\alpha)\in K^{a\#}$, then $K\cap\ell(C_{\psi}(x^{*},\alpha))=\{0\}$.
  • $5^{\circ}$.
    Nontrivialness: If $(x^{*},\alpha)\in K^{a+}$, then $C_{\psi}(x^{*},\alpha)\neq\{0\}$. If $x^{*}\neq 0$ or $\alpha\psi(y)>0$ for some $y\in Y\setminus\{0\}$, then $C_{\psi}(x^{*},\alpha)\neq Y$.
  • $6^{\circ}$.
    Solidness: Assume that $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ (e.g. if either $(x^{*},\alpha)\in K^{a\#}$ or $(x^{*},\alpha)\in K^{a\circ}$ and $K$ is solid). Then, $\operatorname{cor}\,C_{\psi}(x^{*},\alpha)\neq\emptyset,$ and if either $(x^{*},\alpha)\in K^{a\circ}$ and $K$ is solid or $\varphi_{x^{*},\alpha}$ (respectively, $\psi$) is continuous, then $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset.$
  • $7^{\circ}$.
    Lineality: If $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ (e.g. if $(x^{*},\alpha)\in K^{a\#}$), and either $x^{*}\neq 0$ or $\alpha\psi(y)>0$ for some $y\in Y\setminus\{0\}$, then $C_{\psi}(x^{*},\alpha)$ is not a linear subspace (or equivalently, $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$).
  • $8^{\circ}$.
    Dilating cone for $K$: If $(x^{*},\alpha)\in K^{a\#}$, then $C_{\psi}(x^{*},\alpha)$ is a dilating cone for $K$, i.e., $C_{\psi}(x^{*},\alpha)$ is a nontrivial, convex cone with $K\setminus\{0\}\subseteq\operatorname{cor}\,C_{\psi}(x^{*},\alpha)$, and if further $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, then $K\setminus\{0\}\subseteq\operatorname{int}\,C_{\psi}(x^{*},\alpha)$.
Lemma 2.12
For any $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$, the following assertions are valid:
  • $1^{\circ}$.
    $-C_{\psi}(x^{*},\alpha)=S^{\leq}_{\psi}(x^{*},\alpha)$ and $-C^{>}_{\psi}(x^{*},\alpha)=S^{<}_{\psi}(x^{*},\alpha)$.
  • $2^{\circ}$.
    $\varphi_{x^{*},\alpha}$ is sublinear (convex) and $-C_{\psi}(x^{*},\alpha)$ representing.
  • $3^{\circ}$.
    $C^{>}_{\psi}(x^{*},\alpha)\subseteq\operatorname{cor}\,C_{\psi}(x^{*},\alpha)$, and if one has $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$, then $C^{>}_{\psi}(x^{*},\alpha)=\operatorname{cor}\,C_{\psi}(x^{*},\alpha)$.
  • $4^{\circ}$.
    If $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, then $\varphi_{x^{*},\alpha}$ is strictly $-C_{\psi}(x^{*},\alpha)$ representing.
Proof 2.13.
$1^{\circ}$ Both equalities follow from the definitions.
$2^{\circ}$ The sublinearity (convexity) of $\varphi_{x^{*},\alpha}$ follows from the sublinearity (convexity) of both $x^{*}$ and $\psi$. In view of $1^{\circ}$, $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$ representing.
$3^{\circ}$ This assertion follows from Jahn [43, Prop. 3.6] taking into account $1^{\circ}$.
$4^{\circ}$ Of course, $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$ representing by $2^{\circ}$. Assume that $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$. By $1^{\circ}$ and $3^{\circ}$ we get $-S^{<}_{\psi}(x^{*},\alpha)=C^{>}_{\psi}(x^{*},\alpha)=$. Since $C_{\psi}(x^{*},\alpha)$ is a convex set with $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ it follows that $\operatorname{int}\,C_{\psi}(x^{*},\alpha)=$. Thus, we conclude $-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=-$ which shows that $\varphi_{x^{*},\alpha}$ is strictly $-C_{\psi}(x^{*},\alpha)$ representing.
Remark 2.14.
Note that the properties of $S^{\leq}_{\psi}(x^{*},\alpha)$ and $S^{<}_{\psi}(x^{*},\alpha)$, and the relationships between $S^{\leq}_{\psi}(x^{*},\alpha)$ and $S^{<}_{\psi}(x^{*},\alpha)$ follow immediately from the one related to $C_{\psi}(x^{*},\alpha)$ and $C^{>}_{\psi}(x^{*},\alpha)$. Furthermore, note that the conditions $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ (as given in the assumptions of Lemma 2.12 ($4^{\circ}$)) do not imply each other, as can be seen below:
Let us first give an example with $C^{>}_{\psi}(x^{*},\alpha)=\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$. According to [37, Ex. 2.1], for $Y:=\operatorname{\mathbb{R}}^{2}$, $\psi(\cdot):=||\cdot||_{1}$, $x^{*}:=(1,1)$ with $||x^{*}||_{\infty}=1,\alpha:=1$, one has $C_{\psi}(x^{*},1)=\{y\in\operatorname{\mathbb{R}}^{2}\mid x^{*}(y)=||y||_{1}\}=\operatorname{\mathbb{R}}^{2}_{+}$ and in this case $C^{>}_{\psi}(x^{*},\alpha)=\emptyset$, but $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$.
Next, we give an example with $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset$. Recall that, for a set $\Omega\subseteq Y$ in a real normed space $(Y,||\cdot||)$, one denotes the weak interior and weak closure of $\Omega$ by ${\mathrm{i}nt}_{w}\,\Omega$ and ${\mathrm{c}l}_{w}\,\Omega$, respectively. We have the next result:
Proposition 2.15.
Consider a closed, pointed cone $C\subseteq Y$ in a real infinite dimensional normed space $(Y,||\cdot||$). Then, we have $\operatorname{int}_{w}\,C=\emptyset$.
Proof 2.16.
On the contrary, assume that $\operatorname{int}_{w}\,C\neq\emptyset$. Take some $c\in\operatorname{int}_{w}\,C$, i.e., there is a weakly open neighbourhood $U(c)\subseteq Y$ around $c$ which satisfies $U(c)\subseteq C$. Since $Y$ has infinite dimension, the set $U(c)$ contains a line $c+\mathbb{R}\cdot v$ for some $v\in Y\setminus\{0\}$. This implies, for any $n\in\mathbb{N}$, we have $cn^{-1}\pm v=n^{-1}(c\pm nv)\in C$. Taking the limit $n\to\infty$, using the closedness of $C$, we conclude ±$\pm v\in C$, or equivalently, $v\in C\cap(-C)=\ell(C)=\{0\}$, a contradiction.
As a desired second example, we endow a real infinite dimensional (normed) space $Y$ with the weak topology, we put $\psi(\cdot):=||\cdot||$ and we assume $||x^{*}||^{*}>\alpha>0$. In this case, it is known that $C^{>}_{||\cdot||}(x^{*},\alpha)=\operatorname{int}_{||\cdot||}\,C_{||\cdot||}(x^{*},\alpha)\neq\emptyset$. Since $C_{||\cdot||}(x^{*},\alpha)$ is a closed, pointed cone in $(Y,||\cdot||)$ by Lemma 2.11 ($1^{\circ}$, $3^{\circ}$, $4^{\circ}$), we conclude that $\operatorname{int}_{w}\,C_{||\cdot||}(x^{*},\alpha)=\emptyset$ by Proposition 2.15.
Remark also that, if $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\psi$ is continuous, then $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ by Lemma 2.11 ($6^{\circ}$). In the second example from above, the weak upper semicontinuity of |||||$\cdot$||| | ⋅ | | is not satisfied, since $Y$ has infinite dimension (cf. [29, Rem. 2.2]).
2.4 Solution concepts in vector optimization
Given two real topological-linear spaces $X$ and $Y$, a seminorm $\psi:Y\to\mathbb{R}$, a nonempty feasible set $\Omega\subseteq X$, and a vector-valued objective function $f:X\rightarrow Y$, we consider the following vector optimization problem:
$\begin{cases}f(x)\to\min\mbox{ with respect to }K\\ x\in\Omega,\end{cases}$ ($P$)
where we assume throughout the paper that $K\subseteq Y$ is a cone. It is well-known that $K$ induces on $Y$ a reflexive binary relation $\leq_{K}$ defined, for any two points $y,\overline{y}\in Y$, by $y\leq_{K}\overline{y}:\Longleftrightarrow y\in\overline{y}-K.$ Note that if $K$ is convex, then $\leq_{K}$ is a preorder; if $K$ is pointed and convex, then $\leq_{K}$ is a partial order. For notational convenience, we consider further the binary relations $\lneq_{K}$ and $<_{K}$ that are defined, for any two points $y,\overline{y}\in Y$, by $y\lneq_{K}\overline{y}:\Longleftrightarrow y\in\overline{y}-K\setminus\{0\},y<_{K}\overline{y}:\Longleftrightarrow y\in\overline{y}-\operatorname{int}\,K$, respectively. Two well-known types of solutions to the vector problem ($P$) can be defined according to the next two definitions (see, e.g., [5], [23], [42], [49]).
Definition 2.17 (Efficiency).
A point $\overline{x}\in\Omega$ is said to be an efficient solution of ($P$) if there is no $x\in\Omega$ such that $f(x)\lneq_{K}f(\overline{x})$. The set of all Pareto efficient solutions of ($P$) is denoted by ${\mathrm{E}ff}(\Omega\mid f,K)$.
Definition 2.18 (Weak efficiency).
A point $\overline{x}\in\Omega$ is said to be a weakly efficient solution of ($P$) if there is no $x\in\Omega$ such that $f(x)<_{K}f(\overline{x})$. The set of all weakly Pareto efficient solutions of ($P$) is denoted by ${\mathrm{W}Eff}(\Omega\mid f,K)$.
In addition to the efficiency concepts above, we also want to consider a proper efficiency with respect to a cone-valued map.
Definition 2.19 (Proper efficiency with respect to a cone-valued map).
Consider a cone-valued map $\mathbb{A}:\Omega\to 2^{Y}$. A point $\overline{x}\in\Omega$ is said to be an $\mathbb{A}$-properly efficient solution of ($P$) if $\overline{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$ and $\mathbb{A}(\bar{x})\cap(-K)=\{0\}$. The set of all $\mathbb{A}$-properly efficient solutions of ($P$) is denoted by ${\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$.
Remark 2.20.
In Definition 2.19 one can remove the assumption $\overline{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$ if $\mathbb{A}(\bar{x})$ contains $f[\Omega]-f(\bar{x})$. Notice that ${\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)={\mathrm{E}ff}(\Omega\mid f,K)$ and ${\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,\{0\}\cup\operatorname{int}\,K)={\mathrm{W}Eff}(\Omega\mid f,K)$ with $\mathbb{A}(\cdot):=\mathbb{R}_{+}\cdot(f[\Omega]-f(\cdot))$ or $\mathbb{A}(\cdot):=\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\cdot)).$
Remark 2.21.
For the particular choice of the cone-valued map $\operatorname{\mathbb{A}}:\Omega\to 2^{Y}$, one can see that several well-known concepts of proper efficiency are special cases of Definition 2.19: Hurwicz proper efficiency ([40], 1958), Borwein proper efficiency ([3], 1977), Benson proper efficiency ([1], 1979), extended Borwein proper efficiency ([4], 1980). For more details see, e.g., [5, Sec. 2.4.3], [25, Ch. 3], [26], [49, Sec. 2.4] and references therein.
Beside the proper efficiency concept given in Definition 2.19, we like also to consider a proper efficiency concept in the sense of Henig ([38], 1982). To formulate this concept in the next definition, consider the following family of dilating cones:
$\mathcal{D}(K):=\{C\subseteq Y\mid C\text{ is a nontrivial, convex cone with }K\setminus\{0\}\subseteq\operatorname{int}\,C\}.$
Definition 2.22 (Proper efficiency in the sense of Henig).
A point $\bar{x}\in\Omega$ is said to be a Henig properly efficient solution of ($P$) if there is $C\in\mathcal{D}(K)$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C)$. The set of all Henig properly efficient solutions of ($P$) is denoted by $\operatorname{PEff}_{He}(\Omega\mid f,K)$.
The following relationships between the different solution concepts are valid. The proof is straightforward.
Proposition 2.23.
Consider a cone $K\subseteq Y$ with $0\notin\operatorname{int}\,K$ (or equivalently, $K\neq Y$). Then, the following assertions hold:
  • $1^{\circ}$.
    $\operatorname{PEff}_{He}(\Omega\mid f,K)\subseteq\operatorname{Eff}(\Omega\mid f,K)\subseteq\operatorname{WEff}(\Omega\mid f,K)$.
  • $2^{\circ}$.
    $\operatorname{Eff}(\Omega\mid f,C)=\operatorname{WEff}(\Omega\mid f,K)$ for $C:=(\operatorname{int}\,K)\cup\{0\}$.
  • $3^{\circ}$.
    $\operatorname{PEff}_{He}(\Omega\mid f,K)=\bigcup_{C\in\mathcal{D}(K)}\operatorname{Eff}(\Omega\mid f,C)=\bigcup_{C\in\mathcal{D}(K)}\operatorname{WEff}(\Omega\mid f,C)$.
3 Scalarization in Vector Optimization
Under the assumptions of Section 2.4, consider the following scalarization of the vector optimization problem ($P$):
$\begin{cases}(\varphi\circ f)(x)\to\min\\ x\in\Omega,\end{cases}$ ($P_{\varphi}$)
where $\varphi:Y\to\mathbb{R}$ is a (frequently parameter-dependent) function. Clearly, to get useful relationships between the problems ($P$) and ($P_{\varphi}$) one needs to impose certain properties on $\varphi$. By solving the scalar problem ($P_{\varphi}$) (with a specific function $\varphi$), one can also get some knowledge about the original vector problem ($P$). The application of such a strategy is called scalarization method, while the function $\varphi$ is called scalarizing function (see e.g. [5], [23], [42], [49]).
In this paper, we like to analyze the relationships between solutions of ($P_{\varphi}$) and the (weakly, properly) efficient solutions of ($P$) for the case that $\varphi$ satisfies certain monotonicity properties. For doing this, we recall monotonicity concepts for the function $\varphi$ (c.f. Jahn [42, Def. 5.1]).
Definition 3.1.
A scalarizing function $\varphi:Y\to\mathbb{R}$ is said to be
  • $K$-increasing if $\forall\,y,\overline{y}\in Y,\,y\leq_{K}\overline{y}:\;\varphi(y)\leq\varphi(\overline{y}).$
  • strictly $K$-increasing if $\forall\,y,\overline{y}\in Y,\,y<_{K}\overline{y}:\;\varphi(y)<\varphi(\overline{y}).$
  • strongly $K$-increasing if $\forall\,y,\overline{y}\in Y,\,y\lneq_{K}\overline{y}:\;\varphi(y)<\varphi(\overline{y}).$
As an immediate consequence, any strongly $K$-increasing function is $K$-increasing, and if $K\neq Y$, strictly $K$-increasing as well.
Lemma 3.2.
Consider the subadditive functional $\varphi:Y\to\mathbb{R}$ and let $K,C$ be two cones in $Y$. The following assertions hold:
  • $1^{\circ}$.
    If $\varphi$ is (strictly) $-K$ representing, then $\varphi$ is (strictly) $K$-increasing.
  • $2^{\circ}$.
    If $\operatorname{int}K\subseteq\operatorname{int}C$ and $\varphi$ is strictly $C$-increasing, then $\varphi$ is strictly $K$-increasing.
  • $3^{\circ}$.
    If $K\subseteq C$ and $\varphi$ is (strongly) $C$-increasing, then $\varphi$ is (strongly) $K$-increasing.
Proof 3.3.
$1^{\circ}$ We only prove the strict case, since the other case is similar. The case $\operatorname{int}\,K=\emptyset$ is clear. Now assume that $\operatorname{int}\,K\neq\emptyset$. Consider $y^{1},y^{2}\in Y$ with $y^{2}<_{K}y^{1}$, i.e., $y^{2}-y^{1}\in-\operatorname{int}\,K$. Since $\varphi$ is strictly $-K$ representing, this is equivalent to $\varphi(y^{2}-y^{1})<0$. Thus, we derive $\varphi(y^{2})=\varphi(y^{1}+y^{2}-y^{1})\leq\varphi(y^{1})+\varphi(y^{2}-y^{1})<\varphi(y^{1}).$ This shows that $\varphi$ is strictly $K$-increasing. The proofs of $2^{\circ}$ and $3^{\circ}$ are straightforward.
The next scalarization results involving cone-monotone functions are well-known.
Lemma 3.4 (c.f. [50, Lem. 5.6], [27, Lem. 5.2]).
Consider a real-valued function $\varphi:Y\to\mathbb{R}$ and a cone $K\subseteq Y$. Then, the following assertions hold:
  • $1^{\circ}$
    If $\varphi$ is strictly $K$-increasing, then ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi\circ f)(x)\subseteq{\mathrm{W}Eff}(\Omega\mid f,K)$.
  • $2^{\circ}$.
    If $\varphi$ is strongly $K$-increasing, then ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi\circ f)(x)\subseteq{\mathrm{E}ff}(\Omega\mid f,K)$.
  • $3^{\circ}$.
    If $\varphi$ is $K$-increasing, and ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi\circ f)(x)=\{\overline{x}\}$ for some $\overline{x}\in\Omega$, then $\overline{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$.
  • $4^{\circ}$.
    If $\varphi$ is strictly $C$-increasing or strongly $C$-increasing for some $C\in\mathcal{D}(K)$, then ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi\circ f)(x)\subseteq{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$.
For a (strictly) $-K$ representing function $\varphi:Y\to\mathbb{R}$ with $\varphi(0)=0$ we have the following result.
Lemma 3.5 (cf. [43, Th. 4.1]).
Suppose that $K\subseteq Y$ is a cone. Assume that $\varphi:Y\to\mathbb{R}$ is a function with $\varphi(0)=0$. Then, for any $\bar{x}\in\Omega$, we have:
  • $1^{\circ}$.
    If $\varphi$ is $-K$ representing, then
    $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)\iff\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi(f(x)-f(\bar{x})).$
  • $2^{\circ}$.
    If $\varphi$ is strictly $-K$ representing, then
    $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)\iff\bar{x}\in{\mathrm{a}rgmin}_{x\in\Omega}\;\varphi(f(x)-f(\bar{x})).$
Next, we recall two well-known nonlinear scalarization methods in vector optimization which are due to Pascoletti and Serafini [53], and Gerstewitz [20, 21]. For a-priori given parameters $a\in Y$ and $k\in Y\setminus\{0\}$, in [53] the following scalar optimization problem is used
$\begin{cases}\lambda\to\min\\ f(x)\in\lambda k+(a-K),\\ (x,\lambda)\in\Omega\times\mathbb{R},\end{cases}$ (3.1)
while in [20, 21] the scalar optimization problem is given as
$\begin{cases}\varphi^{G}_{a-K,k}(f(x))\to\min\\ x\in\Omega,\end{cases}$ (3.2)
where the nonlinear scalarizing function $\varphi^{G}_{a-K,k}:Y\to\overline{\mathbb{R}}:=\mathbb{R}\cup\{-\infty,+\infty\}$ is defined (for $k\in K\setminus\ell(K)$ and $K$ closed) by
$\varphi^{G}_{a-K,k}(y):=\inf\{t\in\mathbb{R}\mid y\in tk+(a-K)\}\quad\text{forall }y\in Y.$ (3.3)
Many facts about the problems (3.1) and (3.2), as well as about the nonlinear scalarizing function $\varphi^{G}_{a-K,k}$, are known (see, e.g., [20, 21], [53], [22], [12, 13], [49], [57], [23]). For the sake of completeness, we recall some important facts below.
Lemma 3.6.
Suppose that $K\subseteq Y$ is a closed, convex cone and take $k\in K\setminus\ell(K)$. Then, the scalar problems (3.1) and (3.2) are equivalent (i.e., each problem has an optimal solution if and only if the other one has an optimal solution, and the optimal solutions $x$ as well as the optimal value of both problems coincide).
Proof 3.7.
Under the given assumptions, the function $\varphi^{G}_{a-K,k}$ has the level set property $f(x)\in\lambda k+(a-K)\iff\varphi^{G}_{a-K,k}(f(x))\leq\lambda$ (see [23, Th. 2.4.1]). Thus, (3.1) is equivalent to
$\begin{cases}\lambda\to\min\\ \varphi^{G}_{a-K,k}(f(x))\leq\lambda,\\ (x,\lambda)\in\Omega\times\mathbb{R},\end{cases}$
respectively, to (3.2).
In the next lemma, we state classical relationships between the vector problem ($P$) and the scalar problem (3.1) (see, e.g., Eichfelder [12, Th. 2.1 and Cor. 2.9]).
Lemma 3.8.
Suppose that $K\subseteq Y$ is a convex cone. Then, we have:
  • $1^{\circ}$.
    If $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$, then $(\bar{x},0)$ solves the problem (3.1) for $a=f(\bar{x})$ and any $k\in K\setminus\ell(K)$. More precisely, $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the problem (3.1).
  • $2^{\circ}$.
    If $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$, then $(\bar{x},0)$ solves the problem (3.1) for $a=f(\bar{x})$ and any $k\in\operatorname{int}\,K$.
  • $3^{\circ}$.
    Consider $(\bar{x},\bar{\lambda})\in\Omega\times\operatorname{\mathbb{R}}$. If $\{(x,\bar{\lambda})\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the problem (3.1) for some $a\in Y$ and $k\in Y\setminus\{0\}$, then $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$.
  • $4^{\circ}$.
    If $(\bar{x},\bar{\lambda})\in\Omega\times\operatorname{\mathbb{R}}$ solves the problem (3.1) for some $a\in Y$ and $k\in Y\setminus\{0\}$, then $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$ and $f(\bar{x})\in\bar{\lambda}k+(a-\operatorname{bd}\,K)$.
Alternatively, one can immediately obtain scalarization results by using the relationships between (3.1) and (3.2), and properties of the nonlinear scalarizing function $\varphi^{G}_{a-K,k}$ defined in (3.3) (see [57, Ch. 6] for an extended overview on this topic).
Lemma 3.9.
Suppose that $K\subseteq Y$ is a closed, convex cone, and take $k\in K\setminus\ell(K)$. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
  • $1^{\circ}$.
    $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$
  • $2^{\circ}$.
    $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi^{G}_{f(\bar{x})-K,k}(f(x))$.
  • $3^{\circ}$.
    $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the solution set of the problem (3.1) for $a=f(\bar{x})$.
Lemma 3.10.
Suppose that $K\subseteq Y$ is a nontrivial, closed, solid, convex cone, and take $k\in\operatorname{int}\,K$. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
  • $1^{\circ}$.
    $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$
  • $2^{\circ}$.
    $\bar{x}\in{\mathrm{a}rgmin}_{x\in\Omega}\;\varphi^{G}_{f(\bar{x})-K,k}(f(x))$.
  • $3^{\circ}$.
    $(\bar{x},0)$ is a solution of the problem (3.1) for $a=f(\bar{x})$.
The scalar optimization problems of type (3.1) and (3.2) involved in the classical scalarization results given in Lemmata 3.8, 3.9 and 3.10 have the drawback that the parameter $a$ depends on the solution $\bar{x}$ (more precisely, we have $a=f(\bar{x})$). In order to avoid this disadvantage, Jahn [43, Th. 4.2] stated a scalarization result where the parameter $a$ does not depend on the solution $\bar{x}$.
In the following proposition, we present a new scalarization result that contains the result of Jahn [43, Th. 4.2] as a special case (see Remark 3.13 for details).
Proposition 3.11.
Suppose that $K\subseteq Y$ is a convex cone, $\mathcal{P}_{a}\subseteq Y$ is a set of reference points, $\mathcal{P}_{k}\subseteq K\setminus\ell(K)$ is a set of directions and $\mathcal{R}\subseteq\mathbb{R}$ is a set of numbers such that $f[{\mathrm{E}ff}(\Omega\mid f,K)]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
  • $1^{\circ}$.
    $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$.
  • $2^{\circ}$.
    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $\{(x,s)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of (3.1).
If, in addition, $K$ is closed, then $1^{\circ}$ and $2^{\circ}$ are equivalent to:
  • $3^{\circ}$.
    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $\varphi^{G}_{a-K,k}(f(\bar{x}))=s$ and $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi^{G}_{a-K,k}(f(x)).$
Proof 3.12.
The implication $2^{\circ}\Longrightarrow 1^{\circ}$ is a direct consequence of Lemma 3.8 ($3^{\circ}$).
Now, we show the implication $1^{\circ}\Longrightarrow 2^{\circ}$. Let $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$. Take any $k\in\mathcal{P}_{k}\,(\subseteq K\setminus\ell(K))$. By Lemma 3.8 ($1^{\circ}$), the set of solutions of the problem (3.1) (with $f(\bar{x})$ in the role of $a$) is given by $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$. Note that, for any $x\in\Omega$, the pair ($x,0$) solves (3.1) (with $f(\bar{x})$ in the role of $a$) if and only if ($x,t$) solves (3.1) with $f(\bar{x})-tk$ in the role of $a$ for all $t\in\mathbb{R}$. By our assumption $f[{\mathrm{E}ff}(\Omega\mid f,K)]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$, there are $\bar{a}\in\mathcal{P}_{a}$, $\bar{k}\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ such that $f(\bar{x})=\bar{a}+s\bar{k}$. Thus, $\{(x,s)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the problem (3.1) with $\bar{a}=f(\bar{x})-s\bar{k}\in\mathcal{P}_{a}$ and $\bar{k}\in\mathcal{P}_{k}$ in the role of $a$ and $k$.
If $K$ is closed, the equivalence of $2^{\circ}$ and $3^{\circ}$ follows immediately from Lemma 3.6.
Remark 3.13.
[…]
For the concept of weak efficiency we have the following result.
Proposition 3.14.
Suppose that $K\subseteq Y$ is a convex cone, $\mathcal{P}_{a}\subseteq Y$ is a set of reference points, $\mathcal{P}_{k}\subseteq\operatorname{int}\,K$ is a set of directions and $\mathcal{R}\subseteq\mathbb{R}$ is a set of numbers such that $f[{\mathrm{W}Eff}(\Omega\mid f,K)]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
  • $1^{\circ}$.
    $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$.
  • $2^{\circ}$.
    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $(x,s)$ is a solution of (3.1).
If, in addition, $K$ is closed and nontrivial, then $1^{\circ}$ and $2^{\circ}$ are equivalent to:
  • $3^{\circ}$.
    There are some $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $s=\varphi^{G}_{a-K,k}(f(\bar{x}))={\mathrm{m}in}_{x\in\Omega}\;\varphi^{G}_{a-K,k}(f(x)).$
Proof 3.15.
The implication $2^{\circ}\Longrightarrow 1^{\circ}$ is an immediate consequence of Lemma 3.8 ($4^{\circ}$) taking into account that $\operatorname{int}\,K\subseteq K\setminus\{0\}$, or equivalently $K\neq Y$ (since (3.1) has a solution), while the proof of the implication $1^{\circ}\Longrightarrow 2^{\circ}$ is similar to the proof of the corresponding implication in Proposition 3.11.
If $K$ is closed, the equivalence of $2^{\circ}$ and $3^{\circ}$ follows immediately from Lemma 3.6.
Remark 3.16.
Assuming that $K$ has a normlike-base $B_{K}(\psi)$ (which means that the seminorm $\psi:Y\to\mathbb{R}$ is positive on $K\setminus\{0\}$) and $f[{\mathrm{E}ff}(\Omega\mid f,K)]\subseteq a+K\setminus\ell(K)$ (respectively, $f[{\mathrm{W}Eff}(\Omega\mid f,K)]\subseteq a+\operatorname{int}\,K$) for some $a\in Y$, in Proposition 3.11 (respectively, Proposition 3.14), one may take $\mathcal{P}_{a}:=\{a\}$, $\mathcal{P}_{k}:=B_{\{0\}\cup(K\setminus\ell(K))}(\psi)$ (respectively, $\mathcal{P}_{k}:=B_{\{0\}\cup\operatorname{int}\,K}(\psi)$ if $0\notin\operatorname{int}\,K$, or equivalently, $K\neq Y$) and $\mathcal{R}:=\mathbb{P}$.
The following proposition provides possible choices for the sets $\mathcal{P}_{a}\subseteq Y$, $\mathcal{P}_{k}\subseteq K\setminus\ell(K)$ and $\mathcal{R}\subseteq\mathbb{R}$ to ensure $f[D]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$ with $D={\mathrm{E}ff}(\Omega\mid f,K)$, $D={\mathrm{W}Eff}(\Omega\mid f,K)$ and $D=\Omega$, respectively, in Propositions 3.11 and 3.14.
Proposition 3.17.
Let $K\subseteq Y$ be a nontrivial, convex cone and let $y^{*}\in Y^{*}\setminus\{0\}$, $a\in Y$ and $k\in K\setminus\{0\}$ be given. Suppose that $y^{*}(k)>0$ (e.g. if $y^{*}\in K^{\#}$). Define
$H_{y^{*},a}^{\sim}:=\{y\in Y\mid y^{*}(y)\sim y^{*}(a)\}\quad\text{for all }\sim\in\{=,<,>\}.$
Then, the following assertions are valid:
  • $1^{\circ}$.
    $H_{y^{*},a}^{>}=H_{y^{*},a}^{=}+\mathbb{P}\cdot k$ and $H_{y^{*},a}^{<}=H_{y^{*},a}^{=}-\mathbb{P}\cdot k$.
  • $2^{\circ}$.
    If $y^{*}\in K^{\#}$, then $a+K\setminus\{0\}\subseteq H_{y^{*},a}^{=}+\mathbb{P}\cdot k$ and $a-K\setminus\{0\}\subseteq H_{y^{*},a}^{=}-\mathbb{P}\cdot k$.
  • $3^{\circ}$.
    If $y^{*}\in K^{+}$ and $k\in\operatorname{int}\,K$, then $a+\operatorname{int}\,K\subseteq H_{y^{*},a}^{=}+\mathbb{P}\cdot k$ and $a-\operatorname{int}\,K\subseteq H_{y^{*},a}^{=}-\mathbb{P}\cdot k$.
  • $4^{\circ}$.
    $Y=\bigcup_{\sim\in\{<,=,>\}}\,H_{y^{*},a}^{\sim}=H_{y^{*},a}^{=}+\mathbb{R}\cdot k$.
Proof 3.18.
$1^{\circ}$. Let us verify the equality $H_{y^{*},a}^{>}=H_{y^{*},a}^{=}+\mathbb{P}\cdot k$. Since $a\in H_{y^{*},a}^{=}$, $y^{*}(k)>0$ and $\mathbb{P}\cdot k\subseteq K\setminus\{0\}$, we conclude $H_{y^{*},a}^{=}+\mathbb{P}\cdot k\subseteq H_{y^{*},a}^{>}$. Take some $y\in H_{y^{*},a}^{>}$. Noting that $y-tk\in H_{y^{*},a}^{=}$ for some $t>0$, we derive $y=y-tk+tk\in H_{y^{*},a}^{=}+\mathbb{P}\cdot k$, which shows the reverse inclusion. Applying the result for the cone $-K$, the functional $-y^{*}$ and the direction $-k$ (in the role of $K$, $y^{*}$ and $k$), we get immediately $H_{y^{*},a}^{<}=H_{(-y^{*}),a}^{>}=H_{(-y^{*}),a}^{=}+\mathbb{P}\cdot(-k)=H_{y^{*},a}^{=}-\mathbb{P}\cdot k$.
$2^{\circ}$. Assume that $y^{*}\in K^{\#}$, which means that $y^{*}(x)>0$ for all $x\in K\setminus\{0\}$. Then, using $1^{\circ}$ we easily derive $a+K\setminus\{0\}\subseteq H_{y^{*},a}^{>}=H_{y^{*},a}^{=}+\mathbb{P}\cdot k$. Applying the result for $-K$, $-y^{*}$ and $-k$ (in the role of $K$, $y^{*}$ and $k$), we obtain $a-K\setminus\{0\}\subseteq H_{-y^{*},a}^{=}-\mathbb{P}\cdot k=H_{y^{*},a}^{=}-\mathbb{P}\cdot k$.
$3^{\circ}$. Define $C:=({\mathrm{i}nt}\,K)\cup\{0\}$. Note that ${\mathrm{i}nt}\,K=C\setminus\{0\}$ and $K^{+}\setminus\{0\}=\{x^{*}\in Y^{*}\mid\forall x\in{\mathrm{i}nt}\,K:x^{*}(x)>0\}=C^{\#}$ by [50, Th. 4.8, Cor. 4.9] (if $K$ is a solid, convex cone). Applying assertion $2^{\circ}$ for $C$ in the role of $K$ we get the result.
$4^{\circ}$. The first equality is obvious, while the second equality follows by $1^{\circ}$.
Remark 3.19.
Let the assumptions of Proposition 3.17 be fulfilled. Eichfelder’s adaptive scalarization method in multiobjective optimization proposed in [12, 13] is based on parameterized scalar optimization problems of type (3.1) and (3.2), respectively, with different choices of the reference point $a\in H_{y^{*},a}^{=}$ and an a-priori fixed direction $k\in K\setminus\{0\}$. Note that Proposition 3.11 generalizes a result by Eichfelder [13, Th. 3.2] taking into account that $f[{\mathrm{E}ff}(\Omega\mid f,K)]\subseteq Y=\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$ for $\mathcal{P}_{a}:=H_{y^{*},a}^{=}$, $\mathcal{P}_{k}:=\{k\}$ and $\mathcal{R}:=\mathbb{R}$ by Proposition 3.17 ($4^{\circ}$).
4 Nonlinear Cone Separation Conditions
The key tool to prove our main Bishop-Phelps type scalarization results in vector optimization in Section 5.3 will be the separation of two (not necessarily convex) cones by using Bishop-Phelps type separating cones / separating (seminorm-linear) functions. In this section we present all the necessary basics for nonlinear cone separation.
Consider a real topological-linear space $Y$ and a seminorm $\psi:Y\to\operatorname{\mathbb{R}}$. For a cone $C\subseteq Y$, we use the notations:
$S_{C}(\psi):=\operatorname{conv}(B_{C}(\psi))\quad\text{and}\quad S_{C}^{0}(\psi):=\operatorname{conv}(\{0\}\cup B_{C}(\psi)).$
If the underlying seminorm (norm) $\psi$ is clear in the context, we will simply use $S_{C}$ and $S_{C}^{0}$ instead of $S_{C}(\psi)$ and $S_{C}^{0}(\psi)$.
Throughout the next section, we consider a nontrivial cone $K\subseteq Y$ with the normlike-base $B_{K}=B_{K}(\psi)$, and a nontrivial cone $A\subseteq Y$ with the normlike-base $B_{A}=B_{A}(\psi)$. The following conditions will be of interest in our upcoming auxiliary results:
$(\operatorname{cl}\,S_{A}^{0}(\psi))\cap(\operatorname{cl}\,S_{-K}(\psi))=\emptyset,$ (4.1)
$A\cap(\operatorname{cl}(\operatorname{conv}(-K))=\{0\},$ (4.2)
$0\notin\operatorname{cl}\,S_{-K}(\psi).$ (4.3)
The following proposition will later play an important role in proving a Bishop-Phelps type scalarization result (namely Corollary 5.28) for vector optimization problems in real locally convex spaces.
Proposition 4.1.
Assume that (4.2) and (4.3) are valid, and one of the following statements is satisfied:
  1. a)
    $A$ is closed and convex.
  2. b)
    There is $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ such that
    $\operatorname{cl}\,S_{-K}(\psi)$ =$=\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\geq\alpha\}$ (4.4)
    is valid, and one of the following conditions is satisfied:
    $B_{A}\text{ is compact},$ (4.5)
    $S_{A}^{0}\text{ is closed},$ (4.6)
    $A\text{ is closed and }\operatorname{cl}\,B_{A}\text{ is compact},$ (4.7)
    $A\text{ is closed and }\operatorname{conv}(\operatorname{cl}(B_{A})\cup\{0\})\text{ is closed}.$ (4.8)
Then, (4.1) is valid.
Proof 4.2.
a) Assuming $A$ is closed and convex, by (4.2) we get $\operatorname{cl}\,S_{A}^{0}\cap\operatorname{cl}\,S_{-K}\subseteq A\cap\operatorname{cl}(\operatorname{conv}(-K))=\{0\}$, and so by (4.3) we conclude (4.1).
b) Obviously, (4.2) and (4.3) imply $A\cap\operatorname{cl}\,S_{-K}=\emptyset$. Since there is $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ with (4.4), we get that $x^{*}(x)<\alpha$ for all $x\in A\cap(\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi})$. Note that $B_{A}\cup\{0\}\subseteq A\cap\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}$, and if $A$ is closed, then $\operatorname{cl}\,B_{A}\cup\{0\}\subseteq A\cap\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}$. Consider cases:
Case 1: (4.5) is valid. By the continuity of $x^{*}$ and the compactness of $B_{A}\cup\{0\}$, we get some $\beta\in(0,\alpha)$ such that $B_{A}\cup\{0\}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\leq\beta\}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)<\alpha\}$. Hence, $\operatorname{cl}\,S_{A}^{0}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\leq\beta\}$.
Case 2: (4.6) is valid. Then, $\operatorname{cl}\,S_{A}^{0}=S_{A}^{0}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)<\alpha\}$.
Case 3: (4.7) is valid. By the continuity of $x^{*}$ and the compactness of $\operatorname{cl}\,B_{A}\cup\{0\}$ we get some $\beta\in(0,\alpha)$ such that $\operatorname{cl}\,B_{A}\cup\{0\}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\leq\beta\}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)<\alpha\}$. Thus, $\operatorname{cl}\,S_{A}^{0}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\leq\beta\}$.
Case 4: (4.8) is valid. Then, $\operatorname{cl}\,S_{A}^{0}\subseteq\operatorname{cl}(\operatorname{conv}(\operatorname{cl}(B_{A})\cup\{0\}))=\operatorname{conv}(\operatorname{cl}(B_{A})\cup\{0\})\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)<\alpha\}$.
Recalling that (4.4) is valid, we conclude (4.1).
Remark 4.3.
The result in Proposition 4.1 (regarding assumption b)) is a generalization of a result by Kasimbeyli [45, Lem. 3], where the author studied the Euclidean space $Y=\operatorname{\mathbb{R}}^{n}$, a closed cone $A\subseteq Y$ and a closed convex cone $K\subseteq Y$ with $A\cap(-K)=\{0\}$. Moreover, it is assumed that there is $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ such that (4.4) is valid (hence $0\notin\operatorname{cl}\,S_{-K}$ is valid). The conclusion in [45, Lem. 3] is given as the condition (4.1) with $\operatorname{bd}\,A$ in the role of $A$.
The condition (4.4) involved in b) of Proposition 4.1 can be fulfilled by certain Bishop-Phelps type cones, as the following lemma shows.
Lemma 4.4.
Consider a real topological-linear space $Y$, a continuous seminorm $\psi:Y\to\mathbb{R}$, and a nontrivial Bishop-Phelps type cone $C=C_{\psi}(x^{*},\alpha)$ for some $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$. Then,
$\operatorname{cl}\,B_{C}(\psi)=B_{C}(\psi)=\{x\in\operatorname{\mathbb{S}}_{\psi}\mid x^{*}(x)\geq\alpha\},$
and if
$\{e\in Y\setminus\{0\}\mid\psi(e)=1,x^{*}(e)=0\}\neq\emptyset$ (4.9)
(e.g., if $Y$ has dimension two or higher, and $\psi$ is a norm), then
$\operatorname{cl}\,S_{C}(\psi)=S_{C}(\psi)=\{x\in\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\geq\alpha\}.$
Proof 4.5.
The equality $B_{C}=\{x\in\operatorname{\mathbb{S}}_{\psi}\mid x^{*}(x)\geq\alpha\}$ is obvious. Since $\psi$ is continuous, the sets $C$ and $\operatorname{\mathbb{S}}_{\psi}$ are closed, hence $B_{C}$ is closed. Define $D:=\{x\in\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\geq\alpha\}$. Since $D$ is a convex set with $B_{C}\subseteq D$, the inclusion $S_{C}\subseteq D$ is clear.
Let us show that $S_{C}\supseteq D$. Take some $d\in D$. We consider two cases:
Case 1: Assume that $\psi(d)=1$. Then, $d\in B_{C}\subseteq S_{C}$.
Case 2: Assume that $\psi(d)<1$. Put $\gamma:=x^{*}(d)\,(\geq\alpha)$. By our assumption (4.9), there is $e\in Y\setminus\{0\}$ with $\psi(e)=1$ and $x^{*}(e)=0$. Then, $x^{*}(d\pm te)=x^{*}(d)\pm tx^{*}(e)=\gamma$ for all $t\geq 0$, and $\psi(d\pm te)\geq|\,\psi(d)-\psi(te)\,|=|\,\psi(d)-|t|\,|\to\infty$ for $t\to\infty$. Taking into account the continuity of $t\mapsto h_{+}(t):=\psi(d+te)$ and $t\mapsto h_{-}(t):=\psi(d-te)$, as well as the conditions $h_{+}(0)=h_{-}(0)=\psi(d)<1$, $\lim_{t\to\infty}\,h_{+}(t)=\lim_{t\to\infty}\,h_{-}(t)=\infty$, by the well-known intermediate value theorem there are $t_{+},t_{-}\in\operatorname{\mathbb{P}}$ with $h_{+}(t_{+})=h_{-}(t_{-})=1$. Since $x^{*}(d+t_{+}e)=x^{*}(d-t_{-}e)=\gamma\geq\alpha$ we infer $d-t_{-}e\in B_{C}$ and $d+t_{+}e\in B_{C}$, hence $d\in\operatorname{conv}(\{d-t_{-}e,\,d+t_{+}e\})\subseteq S_{C}.$
Of course, $S_{C}=D$ and the closedness of $D$ imply $\operatorname{cl}\,S_{C}=S_{C}$.
Remark 4.6.
Kasimbeyli and Kasimbeyli [44, Th. 4.5] stated a similar result as in Lemma 4.4 in a reflexive Banach space setting (with $\psi(\cdot):=||\cdot||$), but note that the assumption given in (4.9), which in this setting means that $Y$ has dimension two or higher, was not considered in [44, Th. 4.5]. In the one dimensional space ($Y:=\mathbb{R}$, ||$\cdot$|| ⋅ |), the condition (4.9) fails and the conclusion of Lemma 4.4 may fail. To see this, consider $Y:=\mathbb{R}$, $x^{*}:=1$ and $\alpha\in(0,1]$. Then, $C=C(x^{*},\alpha)=\mathbb{R}_{+}$ and $B_{C}(|\cdot|)=\{1\}$ as well as $\operatorname{cl}\,S_{C}(|\cdot|)=\{1\}$ while $\{x\in\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\geq\alpha\}=[\alpha,1]$.
To prove some more specific results in the normed space setting in the upcoming Proposition 4.11 and Corollary 5.30 (see Section 5.3), we need additional lemmata.
Lemma 4.7.
Consider a real normed space $(Y,||\cdot||)$, a (weakly) compact set $\Omega_{1}\subseteq Y$ with $0\notin\Omega_{1}$, and a (weakly) closed, bounded set $\Omega_{2}\subseteq Y$ with $\Omega_{2}\subseteq\operatorname{\mathbb{R}}_{+}\cdot\Omega_{1}$. Then, $\Omega_{2}$ is (weakly) compact.
Proof 4.8.
Note that a set $\Omega_{2}$ in a normed space is weakly bounded if and only if it is (norm) bounded. First, we prove that there exists $t>0$ such that
$\Omega_{2}\subseteq[0,t]\cdot\Omega_{1}.$ (4.10)
On the contrary, assume that there are sequences $(x_{n}^{1})$, $(x_{n}^{2})$ and $(t_{n})$ with $x_{n}^{1}\in\Omega_{1}$, $x_{n}^{2}\in\Omega_{2}$, $t_{n}\in\mathbb{P}$ such that $x_{n}^{2}=t_{n}\cdot x_{n}^{1}$ with $t_{n}\to\infty$ for $n\to\infty$. By the boundedness of $\Omega_{2}$ there is $\lambda>0$ such that $||x||\leq\lambda$ for all $x\in\Omega_{2}$. Consequently, we have $||x_{n}^{1}||=||t_{n}^{-1}x_{n}^{2}||=t_{n}^{-1}||x_{n}^{2}||\leq t_{n}^{-1}\lambda\to 0$ for $n\to\infty$. This means that $0\in\operatorname{cl}\,\Omega_{1}\subseteq\operatorname{cl}_{w}\,\Omega_{1}$. Noting that the compact (respectively, weakly compact) set $\Omega_{1}$ is closed (respectively, weakly closed), we get a contradiction.
Define the function $h:\operatorname{\mathbb{R}}\times Y\to Y$ by $h(s,\omega):=s\cdot\omega$ for all $(s,\omega)\in\operatorname{\mathbb{R}}\times Y$. Since $[0,t]\times\Omega_{1}$ is (weakly) compact in ×$\operatorname{\mathbb{R}}\times Y$, and the function $h$ is (weakly) continuous by the topological vector space properties, we get that $[0,t]\cdot\Omega_{1}$ is (weakly) compact. From (4.10) we derive that the (weakly) closed subset $\Omega_{2}$ of $[0,t]\cdot\Omega_{1}$ is (weakly) compact as well.
Lemma 4.9.
Consider a real normed space $(Y,||\cdot||)$ and a nontrivial cone $C\subseteq Y$. Then, the following assertions are valid:
  • $1^{\circ}$.
    If $B_{C}(||\cdot||)$ is compact, then $B_{C}(||\cdot||^{\prime})$ is compact for all $||\cdot||^{\prime}\sim||\cdot||$.
  • $2^{\circ}$.
    If $\operatorname{cl}_{w}\,B_{C}(||\cdot||)$ is weakly compact with $0\notin\operatorname{cl}_{w}\,B_{C}(||\cdot||)$, then $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ is weakly compact for all $||\cdot||^{\prime}\sim||\cdot||$.
  • $3^{\circ}$.
    If $\operatorname{cl}\,S_{C}(||\cdot||)$ is (weakly) compact with $0\notin\operatorname{cl}\,S_{C}(||\cdot||)$, then $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ is (weakly) compact for all $||\cdot||^{\prime}\sim||\cdot||$.
  • $4^{\circ}$.
    If $Y$ is reflexive (or equivalently, $\operatorname{\mathbb{B}}_{||\cdot||}$ is weakly compact), then $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ and $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ are weakly compact for all $||\cdot||^{\prime}\sim||\cdot||$.
Proof 4.10.
Consider any $||\cdot||^{\prime}\sim||\cdot||$.
$1^{\circ}$. Obviously, $B_{C}(||\cdot||^{\prime})\subseteq C=\operatorname{\mathbb{R}}_{+}\cdot B_{C}(||\cdot||)$. Since $||\cdot||^{\prime}\sim||\cdot||$, there is $s>0$ such that $B_{C}(||\cdot||^{\prime})\subseteq\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\subseteq s\cdot\operatorname{\mathbb{B}}_{||\cdot||}$, which shows that $B_{C}(||\cdot||^{\prime})$ is bounded. Moreover, according to [23, Lemma 2.1.43 (iii)], for the bounded and closed norm-base $B_{C}(||\cdot||)$ with $0\notin B_{C}(||\cdot||)$ we have $C=\mathbb{R}_{+}\cdot B_{C}(||\cdot||)$ is closed, hence $B_{C}(||\cdot||^{\prime})$ is closed as well. Then, the compactness of $B_{C}(||\cdot||^{\prime})$ follows from Lemma 4.7 applied for the compact set $\Omega_{1}:=B_{C}(||\cdot||)$ and the closed, bounded set $\Omega_{2}:=B_{C}(||\cdot||^{\prime})$. Note that $0\notin B_{C}(||\cdot||)=\Omega_{1}$.
$2^{\circ}$. Since $C=\operatorname{\mathbb{R}}_{+}\cdot B_{C}(||\cdot||)\subseteq\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}_{w}\,B_{C}(||\cdot||)\subseteq\operatorname{cl}_{w}\,C$, we have $\operatorname{cl}_{w}\,C=\operatorname{cl}_{w}(\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}_{w}\,B_{C}(||\cdot||))$. Taking into account that $\operatorname{cl}_{w}\,B_{C}(||\cdot||)$ is a weakly closed, bounded set with $0\notin\operatorname{cl}_{w}\,B_{C}(||\cdot||)$, applying [23, Lemma 2.1.43 (iii)] we conclude that $\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}_{w}\,B_{C}(||\cdot||)$ is weakly closed (remark that the weak topology of the normed space $Y$ is Hausdorff). Consequently, $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})\subseteq\operatorname{cl}_{w}\,C=\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}_{w}\,B_{C}(||\cdot||)$. Since $||\cdot||^{\prime}\sim||\cdot||$, there is $s>0$ such that $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})\subseteq\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\subseteq s\cdot\operatorname{\mathbb{B}}_{||\cdot||}$, which shows that $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ is bounded. Then, the weak compactness of $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ follows from Lemma 4.7 applied for the weakly compact set $\Omega_{1}:=\operatorname{cl}_{w}\,B_{C}(||\cdot||)$ and the weakly closed, bounded set $\Omega_{2}:=\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$.
$3^{\circ}$. We have $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})\subseteq\operatorname{cl}(\operatorname{conv}\,C)=\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}\,S_{C}(||\cdot||)$, where the latter equality is discussed in [29, Rem. 2.5]. Since $||\cdot||^{\prime}\sim||\cdot||$, there is $s>0$ such that $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})\subseteq\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\subseteq s\cdot\operatorname{\mathbb{B}}_{||\cdot||}$, hence the set $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ is bounded. Finally, the (weak) compactness of $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ follows from Lemma 4.7 applied for the (weakly) compact set $\Omega_{1}:=\operatorname{cl}\,S_{C}(||\cdot||)$ and the (weakly) closed, bounded set $\Omega_{2}:=\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$.
$4^{\circ}$. Since $||\cdot||^{\prime}\sim||\cdot||$, there is $s>0$ such that $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})\subseteq\operatorname{cl}\,S_{C}(||\cdot||^{\prime})\subseteq\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\subseteq s\cdot\operatorname{\mathbb{B}}_{||\cdot||}$, hence both weakly closed sets $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ and $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ are weakly compact if $\operatorname{\mathbb{B}}_{||\cdot||}$ is weakly compact.
To later prove a Bishop-Phelps type scalarization result (Corollary 5.30) for vector optimization problems in real normed spaces, we need the following proposition.
Proposition 4.11.
Consider a real normed space $(Y,||\cdot||)$ with dimension two or higher, a nontrivial cone $K\subseteq Y$ with the norm-base $B_{K}(||\cdot||)$, a nontrivial cone $A\subseteq Y$ with the norm-base $B_{A}(||\cdot||)$. Suppose that (4.2) and (4.3) with $\psi(\cdot):=||\cdot||$ are valid. Moreover, assume that one of the following statements is valid:
  1. a)
    $B_{A}(||\cdot||)$ is compact.
  2. b)
    $A$ is weakly closed, and either $Y$ is reflexive or $\operatorname{cl}_{w}\,B_{A}(||\cdot||)$ is weakly compact with $0\notin\operatorname{cl}_{w}\,B_{A}(||\cdot||)$.
Then, there is $||\cdot||^{\prime}\sim||\cdot||$ such that (4.1) with $\psi(\cdot):=||\cdot||^{\prime}$ is valid.
Proof 4.12.
Suppose that (4.2) and (4.3) (with $\psi(\cdot):=||\cdot||$) are valid. First, note that the following assertions are known to be equivalent (see [14, Lem. 1.25 and 1.26], [29, Lem. 3.4] and [54, Th. 3.2]):
  • $0\notin\operatorname{cl}\,S_{-K}(||\cdot||)$ (i.e., (4.3) with $\psi(\cdot):=||\cdot||$ is valid);
  • $0\notin\operatorname{cl}\,S_{-\operatorname{cl}(\operatorname{conv}\,K)}(||\cdot||)$;
  • $(-\operatorname{cl}(\operatorname{conv}\,K))^{a+}(||\cdot||)\cap(Y^{*}\times\mathbb{P})\neq\emptyset$;
  • $-\operatorname{cl}(\operatorname{conv}\,K)$ is well-based;
  • $-\operatorname{cl}(\operatorname{conv}\,K)$ is representable as a Bishop-Phelps cone (cf. Example 2.4).
Thus, under assumption (4.3) (with $\psi(\cdot):=||\cdot||$), there is a pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ and an equivalent norm $||\cdot||^{\prime}\sim||\cdot||$ such that $-\operatorname{cl}(\operatorname{conv}\,K)=C_{||\cdot||^{\prime}}(x^{*},\alpha)$. By Lemma 4.4, in the real normed space $(Y,||\cdot||^{\prime})$ with dimension two or higher we have $\operatorname{cl}\,S_{-\operatorname{cl}(\operatorname{conv}\,K)}(||\cdot||^{\prime})=\{x\in\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\mid x^{*}(x)\geq\alpha\}$, hence (4.3) and (4.4) with $\psi(\cdot):=||\cdot||^{\prime}$ and $\operatorname{cl}(\operatorname{conv}\,K)$ in the role of $K$ are valid. Let us finish the proof by assuming a) and b), respectively:
a) In view of Lemma 4.9 ($1^{\circ}$), $B_{A}(||\cdot||^{\prime})$ is compact if $B_{A}(||\cdot||)$ is compact. Proposition 4.1 b) applied for the space $(Y,||\cdot||^{\prime})$, $\psi(\cdot):=||\cdot||^{\prime}$ and $\operatorname{cl}(\operatorname{conv}\,K)$ in the role of $K$ ensures $(\operatorname{cl}\,S_{A}^{0}(||\cdot||^{\prime}))\cap(\operatorname{cl}\,S_{-\operatorname{cl}(\operatorname{conv}\,K)}(||\cdot||^{\prime}))=\emptyset$, hence (4.1) with $\psi(\cdot):=||\cdot||^{\prime}$ is valid, taking into account that $\operatorname{cl}\,S_{{-\mathrm{c}l}(\operatorname{conv}\,K)}(||\cdot||^{\prime})\supseteq\operatorname{cl}\,S_{-K}(||\cdot||^{\prime})$.
b) In view of Lemma 4.9 ($2^{\circ}$, $4^{\circ}$), $\operatorname{cl}_{w}\,B_{A}(||\cdot||^{\prime})$ is weakly compact if either $Y$ is reflexive or $\operatorname{cl}_{w}\,B_{A}(||\cdot||)$ is weakly compact with $0\notin\operatorname{cl}_{w}\,B_{A}(||\cdot||)$. Proposition 4.1 b) applied for the space $Y$ endowed with the weak topology, $\psi(\cdot):=||\cdot||^{\prime}$ and $\operatorname{cl}(\operatorname{conv}\,K)$ in the role of $K$ ensures (4.1) with $\psi(\cdot):=||\cdot||^{\prime}$, taking into account that $\operatorname{cl}_{w}\,S_{A}^{0}(||\cdot||^{\prime})=\operatorname{cl}\,S_{A}^{0}(||\cdot||^{\prime})$ and $\operatorname{cl}_{w}\,S_{-\operatorname{cl}_{w}(\operatorname{conv}\,K)}(||\cdot||^{\prime})=\operatorname{cl}\,S_{{-\mathrm{c}l}(\operatorname{conv}\,K)}(||\cdot||^{\prime})\supseteq\operatorname{cl}\,S_{-K}(||\cdot||^{\prime})$.
Furthermore, we recall some recent results on cone separation in real topological-linear (locally convex) spaces established in [28].
Proposition 4.13 ([28, Th. 5.4]).
Assume that $Y$ is a real topological-linear space, $K,A\subseteq Y$ are nontrivial cones with normlike-bases $B_{K}$ and $B_{A}$, and $K$ is solid. Suppose that $S_{-K}$ is solid and $S_{A}^{0}\cap\operatorname{int}\,S_{-K}=\emptyset.$ Then, there exists $(x^{*},\alpha)\in K^{a\circ}$ such that
$x^{*}(a)+\alpha\psi(a)$ $\geq 0\geq x^{*}(k)+\alpha\psi(k)\quad\mbox{for all }a\in A\mbox{and }k\in-K,$
$x^{*}(a)+\alpha\psi(a)$ $\geq 0>x^{*}(k)+\alpha\psi(k)\quad\mbox{for all }a\in A\mbox{ and}k\in\operatorname{int}(-K),$
or equivalently,
$A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset\quad\text{and}\quad K\subseteq C_{\psi}(x^{*},\alpha)\quad$ (4.11)
$\text{and}\quad\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha).$
Proposition 4.14 ([28, Th. 5.8]).
Assume that $Y$ is a real topological-linear space, $K,A\subseteq Y$ are nontrivial, convex cones with normlike-bases $B_{K}$ and $B_{A}$, and $K$ is solid. Suppose that $S_{-K}$ is solid. Then, the following assertions are equivalent:
  • $1^{\circ}$.
    $S_{A}^{0}\cap\operatorname{int}\,S_{-K}=\emptyset$.
  • $2^{\circ}$.
    $A\cap\operatorname{int}(-K)=\emptyset$.
  • $3^{\circ}$.
    There exists $(x^{*},\alpha)\in K^{a\circ}$ such that (4.11) is valid.
Proposition 4.15 ([28, Th. 5.9]).
Assume that $Y$ is a real locally convex space, $K,A\subseteq Y$ are nontrivial cones with normlike-bases $B_{K}$ and $B_{A}$. Suppose that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{A}^{0}$ is compact, and (4.1) is valid. Then, there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
$x^{*}(a)+\alpha\psi(a)>0>x^{*}(k)+\alpha\psi(k)\quad\mbox{for all }a\in A\setminus\{0\}\mbox{ and }k\in-K\setminus\{0\},$
or equivalently,
$A\cap-C_{\psi}(x^{*},\alpha)=\{0\}\quad\text{and}\quad K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=$ (4.12)
From the general strict cone separation result in Proposition 4.15 and our Propositions 4.1 and 4.11 we immediately derive some new variants of strict cone separation results involving the classical separation condition (4.2).
Corollary 4.16.
Assume that $Y$ is a real locally convex space, $K,A\subseteq Y$ are nontrivial cones with normlike-bases $B_{K}$ and $B_{A}$. Suppose that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{A}^{0}$ is compact, and (4.2) and (4.3) are valid. Moreover, assume that the assumptions in a) or in b) of Proposition 4.1 are satisfied. Then, there is $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that (LABEL:sep_conditions_strict) is valid.
Corollary 4.17.
Consider a real normed space $(Y,||\cdot||)$ with dimension two or higher, a nontrivial cone $K\subseteq Y$ with the norm-base $B_{K}(||\cdot||)$, a nontrivial cone $A\subseteq Y$ with the norm-base $B_{A}(||\cdot||)$. Suppose that one of the sets ${\mathrm{c}l}\,S_{-K}(||\cdot||)$ and ${\mathrm{c}l}\,S_{A}^{0}(||\cdot||)$ is weakly compact (e.g., if $Y$ is reflexive), and (4.2) and (4.3) with $\psi(\cdot):=||\cdot||$ are valid. Moreover, assume that the assumptions in a) or in b) of Proposition 4.11 are satisfied. Then, there are $||\cdot||^{\prime}\sim||\cdot||$ and $(x^{*},\alpha)\in K^{a\#}(||\cdot||^{\prime})\cap(Y^{*}\times\mathbb{P})$ such that (LABEL:sep_conditions_strict) with $\psi(\cdot):=||\cdot||^{\prime}$ is valid.
5 Bishop-Phelps Type Scalarization in Vector Optimization
Assume that the real topological-linear space $Y$ is endowed with a nontrivial cone $K\subseteq Y$. Then, for any pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$, and any given reference point $a\in Y$, one defines $\varphi_{x^{*},\alpha}^{a}:Y\to\operatorname{\mathbb{R}}$ as $\varphi_{x^{*},\alpha}^{a}(y):=\varphi_{x^{*},\alpha}(y-a)$ for all $y\in Y$. The following parameterized scalar optimization problem of the vector problem ($P$) (in the sense of Kasimbeyli [45]) will be of interest:
$\begin{cases}(\varphi_{x^{*},\alpha}^{a}\circ f)(x)=\varphi_{x^{*},\alpha}(f(x)-a)=x^{*}(f(x)-a)+\alpha\psi(f(x)-a)\to\min\\ x\in\Omega.\end{cases}$ ($P_{\varphi}^{a}$)
In addition, for any pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$ and any given reference point $a\in Y$ and direction $k\in Y\setminus\{0\}$, we are interested in the parameterized scalar problem (in the sense of Gerstewitz [20, 21]; Jahn [43]; Pascoletti and Serafini [53]):
$\begin{cases}\lambda\to\min\\ \varphi_{x^{*},\alpha}(f(x)-a-\lambda k)\leq 0,\\ (x,\lambda)\in\Omega\times\mathbb{R}.\end{cases}$ ($P_{\varphi}^{a,k}$)
5.1 Basic scalarization results
First, we present monotonicity properties of the seminorm-linear function $\varphi_{x^{*},\alpha}$ with respect to a Bishop-Phelps type cone $C_{\psi}(x^{*},\alpha)$.
Lemma 5.1.
The following assertions hold:
  • $1^{\circ}$.
    If $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$, then $\varphi_{x^{*},\alpha}$ is $C_{\psi}(x^{*},\alpha)$-increasing.
  • $2^{\circ}$.
    If $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$ and $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$, then $\varphi_{x^{*},\alpha}$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing.
  • $3^{\circ}$.
    If $K$ is solid and $(x^{*},\alpha)\in K^{a\circ}$, then $\varphi_{x^{*},\alpha}$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing.
  • $4^{\circ}$.
    If $(x^{*},\alpha)\in K^{a\#}$, then $\varphi_{x^{*},\alpha}$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing.
  • $5^{\circ}$.
    If $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$, then $\varphi_{x^{*},\alpha}$ is strongly $(C_{\psi}^{>}(x^{*},\alpha)\cup\{0\})$-increasing.
Proof 5.2.
$1^{\circ}$ Noting that $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$ representing, the assertion follows from Lemma 3.2 ($1^{\circ}$).
$2^{\circ}$ If $\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset$, there is nothing to prove. If $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, using the assumption $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ we have that $\varphi_{x^{*},\alpha}$ is strictly $-C_{\psi}(x^{*},\alpha)$ representing, and the result follows from Lemma 3.2 ($1^{\circ}$).
$3^{\circ}$ and $4^{\circ}$ Using the assumptions it follows that $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$, so by assertion $2^{\circ}$ we get the conclusion.
$5^{\circ}$ It follows from the definitions.
The next lemma, which studies the monotonicity properties of $\varphi_{x^{*},\alpha}$ with respect to the original cone $K$, extends a corresponding result derived by Kasimbeyli [45, Th. 3.5] in a normed setting.
Lemma 5.3.
Consider $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$. The following assertions hold:
  • $1^{\circ}$.
    $(x^{*},\alpha)\in K^{a+}$ $\iff$ $\varphi_{x^{*},\alpha}$ is $K$-increasing.
  • $2^{\circ}$.
    $(x^{*},\alpha)\in K^{a\circ}$ $\iff$ $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing.
  • $3^{\circ}$.
    $(x^{*},\alpha)\in K^{a\#}$ $\iff$ $\varphi_{x^{*},\alpha}$ is strongly $K$-increasing.
Proof 5.4.
$1^{\circ}$ By Lemma 5.1 ($1^{\circ}$), the function $\varphi_{x^{*},\alpha}$ is $C_{\psi}(x^{*},\alpha)$-increasing. If $(x^{*},\alpha)\in K^{a+}$, then $K\subseteq C_{\psi}(x^{*},\alpha)$, hence $\varphi_{x^{*},\alpha}$ is $K$-increasing by Lemma 3.2 ($3^{\circ}$). Conversely, assuming $\varphi_{x^{*},\alpha}$ is $K$-increasing, for $k\in K$ we have $-k\leq_{K}0$, hence $\varphi_{x^{*},\alpha}(-k)\leq\varphi_{x^{*},\alpha}(0)=0$, which means that $x^{*}(k)-\alpha\psi(k)\geq 0$. Noting that $x^{*}\in K^{+}$, we conclude $(x^{*},\alpha)\in K^{a+}$.
$2^{\circ}$ If $\operatorname{int}\,K=\emptyset$, we have $K^{a\circ}=Y^{*}\times\mathbb{R}_{+}$ and $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing, hence the equivalence is valid. Assume that $\operatorname{int}\,K\neq\emptyset$. If $(x^{*},\alpha)\in K^{a\circ}$, then $\operatorname{int}\,K\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, and so $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing by Lemmata 5.1 ($2^{\circ}$) and 3.2 ($2^{\circ}$). The proof of the reverse implication is similar to $1^{\circ}$.
$3^{\circ}$ If $(x^{*},\alpha)\in K^{a\#}$, then $K\subseteq C_{\psi}^{>}(x^{*},\alpha)\cup\{0\}$, so $\varphi_{x^{*},\alpha}$ is strongly $K$-increasing by Lemmata 5.1 ($5^{\circ}$) and 3.2 ($3^{\circ}$). Also here the proof of the reverse implication is similar to $1^{\circ}$.
Now, we are able to present first relationships between the parameterized scalar problem ($P_{\varphi}^{a}$) and the vector problem ($P$).
Theorem 5.5.
For any $a\in Y$, the following assertions hold:
  • $1^{\circ}$.
    If $(x^{*},\alpha)\in K^{a\circ}$, then ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi_{x^{*},\alpha}^{a}\circ f)(x)\subseteq{\mathrm{W}Eff}(\Omega\mid f,K)$.
  • $2^{\circ}$.
    If $(x^{*},\alpha)\in K^{a\#}$, then ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi_{x^{*},\alpha}^{a}\circ f)(x)\subseteq{\mathrm{E}ff}(\Omega\mid f,K)$.
  • $3^{\circ}$.
    Assume that $\overline{x}\in\Omega$. If $(x^{*},\alpha)\in K^{a+}$ and ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi_{x^{*},\alpha}^{a}\circ f)(x)=\{\overline{x}\}$, then $\overline{x}\in{\mathrm{E}ff}(\Omega\mid f,K)$.
  • $4^{\circ}$.
    If $(x^{*},\alpha)\in K^{a\#}$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, then ${\mathrm{a}rgmin}_{x\in\Omega}\;(\varphi_{x^{*},\alpha}^{a}\circ f)(x)\subseteq{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$.
Proof 5.6.
$1^{\circ}$ Consider $(x^{*},\alpha)\in K^{a\circ}$. By Lemma 5.3 ($2^{\circ}$) $\varphi_{x^{*},\alpha}$ is strictly $K$-increasing. It is easy to check that $\varphi_{x^{*},\alpha}^{a}$ is strictly $K$-increasing for any $a\in Y$ as well. Finally, Lemma 3.4 ($1^{\circ}$) yields the conclusion in assertion $1^{\circ}$.
$2^{\circ}$ and $3^{\circ}$ The proof of the assertions is similar to the proof of assertion $1^{\circ}$ by using Lemmata 5.3 and 3.4.
$4^{\circ}$ Take some $(x^{*},\alpha)\in K^{a\#}$. By Lemma 5.1 ($4^{\circ}$) $\varphi_{x^{*},\alpha}$ is strictly $C_{\psi}(x^{*},\alpha)$-increasing, and so $\varphi_{x^{*},\alpha}^{a}$ as well. Because $(x^{*},\alpha)\in K^{a\#}$ we get $K\setminus\{0\}\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$. Taking into account that $C_{\psi}(x^{*},\alpha)$ is a nontrivial convex cone, we conclude that $C_{\psi}(x^{*},\alpha)\in\mathcal{D}(K)$. Finally, with the aid of Lemma 3.4 ($4^{\circ}$) we get the desired inclusion.
Remark 5.7.
Consider $\bar{x}\in\Omega$. In some upcoming results, the specific choice $a:=f(\bar{x})$ will be of interest. In this case, the scalarized objective function $\varphi_{x^{*},\alpha}^{a}\circ f$ is given by $(\varphi_{x^{*},\alpha}^{a}\circ f)(x)=\varphi_{x^{*},\alpha}(f(x)-f(\bar{x}))$ for all $x\in\Omega$. According to Theorem 5.5, for a specific choice of the pair $(x^{*},\alpha)$ one can ensure the (weak, proper) efficiency of $\bar{x}\in{\mathrm{a}rgmin}_{x\in\Omega}\;\varphi_{x^{*},\alpha}(f(x)-f(\bar{x}))$.
Let us shift our focus to relationships between the parameterized scalar problem ($P_{\varphi}^{a,k}$) and the vector problem ($P$) involving a Bishop-Phelps type cone.
Lemma 5.8.
Consider $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$. The following assertions are valid:
  • $1^{\circ}$.
    If $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$, then $(\bar{x},0)$ solves the problem ($P_{\varphi}^{a,k}$) for $a=f(\bar{x})$ and any $k\in C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$. More precisely, $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of ($P_{\varphi}^{a,k}$).
  • $2^{\circ}$.
    If $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$, then $(\bar{x},0)$ solves the problem ($P_{\varphi}^{a,k}$) for $a=f(\bar{x})$ and any $k\in\operatorname{int}\,C_{\psi}(x^{*},\alpha)$.
  • $3^{\circ}$.
    Consider $(\bar{x},\bar{\lambda})\in\Omega\times\operatorname{\mathbb{R}}$. If $\{(x,\bar{\lambda})\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of the problem ($P_{\varphi}^{a,k}$) for some $a\in Y$ and $k\in Y\setminus\{0\}$, then $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
  • $4^{\circ}$.
    If $(\bar{x},\bar{\lambda})$ solves the problem ($P_{\varphi}^{a,k}$) for some $a\in Y$ and $k\in Y\setminus\{0\}$, then $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
Proof 5.9.
Follows immediately from Lemma 3.8 taking into account that $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$ representing.
Theorem 5.10.
Consider $\bar{x}\in\Omega$. Take some $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{R}}_{+}$, $a\in Y,k\in Y\setminus\{0\}$ and assume that $(\bar{x},\bar{\lambda})$ is a solution of the optimization problem ($P_{\varphi}^{a,k}$). Then:
  • $1^{\circ}$.
    If $(x^{*},\alpha)\in K^{a\#}$ and $C_{\psi}(x^{*},\alpha)$ is solid, then $\bar{x}\in{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$.
  • $2^{\circ}$.
    If $(x^{*},\alpha)\in K^{a\circ}$, then $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$.
Proof 5.11.
$1^{\circ}$. Remark that $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ in view of Lemma 5.8 ($4^{\circ}$). The latter set is contained in ${\mathrm{P}Eff}_{He}(\Omega\mid f,K)$ (since $C_{\psi}(x^{*},\alpha)\in\mathcal{D}(K)$).
$2^{\circ}$. For the case $\operatorname{int}\,K=\emptyset$ the statement is clear. Assume that $\operatorname{int}\,K\neq\emptyset$. Again we have $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Here the latter set is contained in ${\mathrm{W}Eff}(\Omega\mid f,K)$ (since $\emptyset\neq\operatorname{int}\,K\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$).
Let us summarize our findings related to vector problems involving a Bishop-Phelps type cone in the next two propositions.
Proposition 5.12.
Let $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$ such that $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
  • $1^{\circ}$.
    $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
  • $2^{\circ}$.
    For any $k\in C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$, the set $\{(x,0)\mid x\in\Omega,f(x)=f(\bar{x})\}$ is the set of solutions of ($P_{\varphi}^{a,k}$) for $a=f(\bar{x})$.
  • $3^{\circ}$.
    $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi_{x^{*},\alpha}(f(x)-f(\bar{x}))$.
If there are sets $\mathcal{P}_{a}\subseteq Y$, $\mathcal{P}_{k}\subseteq C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$ and $\mathcal{R}\subseteq\mathbb{R}$ such that $f[{\mathrm{E}ff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$, then $1^{\circ}-3^{\circ}$ are equivalent to:
If further $C_{\psi}(x^{*},\alpha)$ is closed, then $1^{\circ}-4^{\circ}$ are equivalent to:
  • $5^{\circ}$.
    There are $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(\bar{x}))=s$ and $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(x)).$
Proof 5.13.
Note that $C_{\psi}(x^{*},\alpha)$ is a convex cone, $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$-representing with $\varphi_{x^{*},\alpha}(0)=0$ and $C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))\neq\emptyset$.
$1^{\circ}\iff 2^{\circ}$. The equivalence follows immediately from Lemma 5.8 ($1^{\circ}$, $3^{\circ}$).
$1^{\circ}\iff 3^{\circ}$. The result follows directly from Lemma 3.5 ($1^{\circ}$).
$1^{\circ}\iff 4^{\circ}\iff 5^{\circ}$. The result follows from Proposition 3.11 taking into account that ($P_{\varphi}^{a,k}$) is the same problem as (3.1) with $C_{\psi}(x^{*},\alpha)$ in the role of $K$.
Proposition 5.14.
Consider $(x^{*},\alpha)\in Y^{*}\times\mathbb{R}_{+}$ such that $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$. Then, for any $\bar{x}\in\Omega$, the following assertions are equivalent:
  • $1^{\circ}$.
    $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$.
  • $2^{\circ}$.
    For any $k\in\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, the point $(\bar{x},0)$ is a solution of the problem ($P_{\varphi}^{a,k}$) for $a=f(\bar{x})$.
If $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$, then $1^{\circ}$ and $2^{\circ}$ are equivalent to:
  • $3^{\circ}$.
    $\bar{x}\in{\mathrm{a}rgmin}_{x\in\Omega}\;\varphi_{x^{*},\alpha}(f(x)-f(\bar{x}))$.
If there are sets $\mathcal{P}_{a}\subseteq Y$, $\mathcal{P}_{k}\subseteq\operatorname{int}\,C_{\psi}(x^{*},\alpha)$ and $\mathcal{R}\subseteq\mathbb{R}$ such that $f[{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))]\subseteq\mathcal{P}_{a}+\mathcal{R}\cdot\mathcal{P}_{k}$, then $1^{\circ}$ and $2^{\circ}$ are equivalent to:
If further $C_{\psi}(x^{*},\alpha)$ is closed and nontrivial, then $1^{\circ},2^{\circ}$ and $4^{\circ}$ are equivalent to:
  • $5^{\circ}$.
    There are $a\in\mathcal{P}_{a}$, $k\in\mathcal{P}_{k}$ and $s\in\mathcal{R}$ so that $s=\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(\bar{x}))={\mathrm{m}in}_{x\in\Omega}\;\varphi^{G}_{a-C_{\psi}(x^{*},\alpha),k}(f(x)).$
Proof 5.15.
First, note that $C_{\psi}(x^{*},\alpha)$ is a solid, convex cone and $\varphi_{x^{*},\alpha}$ is $-C_{\psi}(x^{*},\alpha)$-representing with $\varphi_{x^{*},\alpha}(0)=0$.
$1^{\circ}\iff 2^{\circ}$. The equivalence follows immediately from Lemma 5.8 ($2^{\circ}$, $4^{\circ}$).
$1^{\circ}\iff 3^{\circ}$. Since $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$ and $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$, the function $\varphi_{x^{*},\alpha}$ is strictly $-C_{\psi}(x^{*},\alpha)$-representing by Lemma 2.12 ($4^{\circ}$) with $\varphi_{x^{*},\alpha}(0)=0$. Then, the result follows from Lemma 3.5 ($2^{\circ}$).
$1^{\circ}\iff 4^{\circ}\iff 5^{\circ}$. The result follows directly from the Proposition 3.14.
Remark 5.16.
The results derived in Theorem 5.5 and Propositions 5.12 and 5.14 extend some known results in the literature, which are given in a normed space $(Y,\psi)$, $\psi(\cdot):=||\cdot||$, preordered by a convex cone $K\subseteq Y$:
  • Gasimov [17, Th. 1] proved a corresponding statement as in Theorem 5.5 ($4^{\circ}$) for Benson proper efficiency (instead of Henig proper efficiency).
  • Kasimbeyli [46, Th. 5.4] stated similar assertions as in Theorem 5.5 for a closed, pointed, convex cone $K$ and Benson proper efficiency (instead of Henig proper efficiency).
  • Gasimov [17, Th. 2] (see also Guo and Zhang [31, Th. 4.5]) stated that $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,C_{||\cdot||}(x^{*},\alpha))$ (with $\mathbb{A}$ considered in Benson proper efficiency) implies $\{x\in\Omega\mid f(x)=f(\bar{x})\}={\mathrm{a}rgmin}_{x\in\Omega}\;\varphi_{x^{*},\alpha}(f(x)-f(\bar{x}))$.
Remark 5.17.
It is worth noting that in Proposition 5.12 (respectively, in Proposition 5.14) the function $\varphi_{x^{*},\alpha}$, which appears in the scalar problems ($P_{\varphi}^{a}$) and ($P_{\varphi}^{a,k}$), can be replaced by any $-C_{\psi}(x^{*},\alpha)$ representing (respectively, strictly $-C_{\psi}(x^{*},\alpha)$ representing) function $\varphi:Y\to\mathbb{R}$ with $\varphi(0)=0$. For instance, the nonlinear scalarizing function $\varphi^{G}_{-C_{\psi}(x^{*},\alpha),k}$ (i.e., the function defined in (3.3) with $a=0$ and with $C_{\psi}(x^{*},\alpha)$ in the role of $K$) is $-C_{\psi}(x^{*},\alpha)$ representing (respectively, strictly $-C_{\psi}(x^{*},\alpha)$ representing) if $C_{\psi}(x^{*},\alpha)$ is nontrivial and closed, and $k\in C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha)$ (respectively, $k\in\operatorname{int}\,C_{\psi}(x^{*},\alpha)$).
Remark 5.18.
Assume that $K$ is a nontrivial cone and $(x^{*},\alpha)\in K^{a\#}$. Then, $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)\subseteq C_{\psi}(x^{*},\alpha)\setminus\{0\}$, hence $\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\mathrm{E}ff}(\Omega\mid f,K)$. If further $K$ has the normlike-base $B_{K}(\psi)$, and $C_{\psi}(x^{*},\alpha)$ is solid (e.g. if $K$ is solid or $\psi$ is continuous), then $C_{\psi}(x^{*},\alpha)$ is a nontrivial, convex cone with $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, which means that $C_{\psi}(x^{*},\alpha)\in\mathcal{D}(K)$, and so (see Proposition 2.23) $\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$.
If $K$ is a nontrivial, solid cone and $(x^{*},\alpha)\in K^{a\circ}$, then $\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$, hence $\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))\subseteq{\mathrm{W}Eff}(\Omega\mid f,K)$.
5.2 Scalarization results for the concept of weak efficiency
In this section, we focus on the solution concept of weak efficiency. More precisely, we will present conditions that ensure that a weakly efficient solution of the vector problem ($P$) is a solution of the scalar problems ($P_{\varphi}^{a}$) and ($P_{\varphi}^{a,k}$) for some parameters $a\in Y$ and $k\in C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$.
Theorem 5.19.
Assume that $Y$ is a real topological-linear space, $K$ is a nontrivial, solid cone, $\bar{x}\in\Omega$, and $A:=\mathbb{A}(\bar{x})=\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\bar{x}))$ is nontrivial. Moreover, suppose that $B_{K}$ and $B_{A}$ are normlike-bases for $K$ and $A$, respectively. Suppose $S_{-K}$ is solid and
$S_{A}^{0}\cap\operatorname{int}\,S_{-K}=\emptyset.$ (5.1)
Then, there is $(x^{*},\alpha)\in K^{a\circ}$ such that $\bar{x}\in\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Moreover, any of the equivalent assertions given in Proposition 5.14 is valid.
Proof 5.20.
Under our assumptions, by Proposition 4.13 there exists $(x^{*},\alpha)\in K^{a\circ}$ such that $A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset$ and $\operatorname{int}\,K\subseteq C_{\psi}^{>}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$. Taking into account that $f[\Omega]-f(\bar{x})\subseteq A$, we have $(f[\Omega]-f(\bar{x}))\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)\subseteq A\cap-\operatorname{int}\,C_{\psi}(x^{*},\alpha)=\emptyset,$ hence $\bar{x}\in\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Since $C^{>}_{\psi}(x^{*},\alpha)\neq\emptyset$ and $\operatorname{int}\,C_{\psi}(x^{*},\alpha)\neq\emptyset$, Proposition 5.14 yields the rest of the conclusion.
Remark 5.21.
In view of Remark 5.18, the point $\bar{x}\in\operatorname{WEff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ given in the conclusion of Theorem 5.19 belongs to ${\mathrm{W}Eff}(\Omega\mid f,K)$.
In the convex case (i.e., $K$ and $\mathbb{A}(\bar{x})$ are convex cones), using Proposition 4.14 we are able to replace the separation condition (5.1) by $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$, as shown in the following result.
Corollary 5.22 (convex case).
Assume that $Y$ is a real topological-linear space, $K$ is a nontrivial, solid, convex cone, $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})=\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\bar{x}))$ is nontrivial and convex. Moreover, suppose that $B_{K}$ and $B_{A}$ are normlike-bases for $K$ and $A$, respectively, and $S_{K}$ is solid.
Then, the conclusion of Theorem 5.19 holds.
Proof 5.23.
Follows from Theorem 5.19 and Proposition 4.14.
Remark 5.24.
Recall that, for any set $\Omega\subseteq Y,$ the quasi-interior of $\Omega$ is the set ${\mathrm{q}i}\,\Omega:=\{x\in\Omega\mid\operatorname{cl}(\mathbb{R}_{+}\cdot(\Omega-x))=Y\}$. In order to ensure that the assumption $A$ is nontrivial and convex from Corollary 5.22 is satisfied it is sufficient to impose that $Y$ is a real Hausdorff locally convex space and $f$ is $K$-convexlike (i.e., $f[\Omega]+K$ is convex). Indeed, assume that $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,K)$. Since $f$ is $K$-convexlike, we infer that $A=\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\bar{x}))$ is a convex cone in $Y$. By [27, Lem. 3.3 (4)] we know that $f(\bar{x})\notin\operatorname{int}(f[\Omega]+K)$. Since the interior and the quasi interior of a solid, convex set $f[\Omega]+K$ in a real Hausdorff locally convex space coincide (see [23, p. 66-67]), we have $A\subseteq{\mathrm{c}l}(\mathbb{R}_{+}\cdot(f[\Omega]+K-f(\bar{x})))\neq Y$. Moreover, since $K\neq\{0\}$ we have $A\neq\{0\}$ as well, hence $A$ is nontrivial.
5.3 Scalarization results for the concept of proper efficiency with respect to a cone-valued map
We are ready to present some further main scalarization results which ensure that an $\mathbb{A}$-properly efficient solution of the vector problem ($P$) is a solution of the scalar problems ($P_{\varphi}^{a}$) and ($P_{\varphi}^{a,k}$) for some parameters $a\in Y$ and $k\in C_{\psi}(x^{*},\alpha)\setminus\ell(C_{\psi}(x^{*},\alpha))$.
Theorem 5.25.
Assume that $Y$ is a real locally convex space, $K\subseteq Y$ is a nontrivial cone, $\bar{x}\in\Omega$, and $A:=\mathbb{A}(\bar{x})$ is a nontrivial cone in $Y$ with $f[\Omega]-f(\bar{x})\subseteq A$. Moreover, suppose that $B_{K}$ and $B_{A}$ are normlike-bases for $K$ and $A$, respectively. Suppose that one of the sets $\operatorname{cl}\,S_{-K}$ and $\operatorname{cl}\,S_{A}^{0}$ is compact, and
$({\mathrm{c}l}\,S_{A}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset.$ (5.2)
Then, there is $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\operatorname{\mathbb{P}})$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Moreover, any of the equivalent assertions given in Proposition 5.12 is valid.
Proof 5.26.
By Proposition 4.15 there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\operatorname{\mathbb{P}})$ such that $A\cap-C_{\psi}(x^{*},\alpha)=\{0\}$ and $K\setminus\{0\}\subseteq C_{\psi}^{>}(x^{*},\alpha)=$. Then, taking into account that $f[\Omega]-f(\bar{x})\subseteq A$, we have $(f[\Omega]-f(\bar{x}))\cap-C_{\psi}(x^{*},\alpha)\subseteq A\cap-C_{\psi}(x^{*},\alpha)=\{0\},$ hence $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Finally, observing that $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$ by Lemma 2.11 ($7^{\circ}$), Proposition 5.12 yields the rest of the conclusion.
Remark 5.27.
In view of Remark 5.18, if $K$ is solid or $\psi$ is continuous, then the point $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ given in the conclusion of Theorem 5.25 belongs to ${\mathrm{P}Eff}_{He}(\Omega\mid f,K)$.
Using Proposition 4.1 we are able to replace the separation condition (5.2) by $0\notin\operatorname{cl}\,S_{-K}$ and $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$ as well as some more specific conditions that we formalize below.
Corollary 5.28.
Assume that $Y$ is a real locally convex space, $K\subseteq Y$ is a nontrivial, closed, convex cone with $0\notin\operatorname{cl}\,S_{-K}$, $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})$ is a cone in $Y$ with $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$. Suppose that $B_{K}$ and $B_{A}$ are normlike-bases for $K$ and $A$, and that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{A}^{0}$ is compact. Moreover, assume that one of the following statements is valid:
  1. a)
    $A$ is closed and convex.
  2. b)
    There is $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ such that (4.4) is valid, and one of the conditions (4.5)–(4.8) holds.
Then, the conclusion of Theorem 5.25 is valid.
Proof 5.29.
Using Proposition 4.1 a) and b) respectively, under our standing assumptions a) and b), we get that $A\cap\operatorname{cl}(\operatorname{conv}(-K))=A\cap(-K)=\{0\}$ and $0\notin\operatorname{cl}\,S_{-K}$ imply (5.2). Note that the nontrivialness of $A$ follows from $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$, $K$ is nontrivial and $A\cap(-K)=\{0\}$. Then, the result follows from Theorem 5.25.
In the remaining part of the section, we present some more particularized results for the case of a real normed space $Y$ by using Proposition 4.11 and Lemma 4.9.
Corollary 5.30.
Assume that ($Y,||\cdot||$) is a real normed space with dimension two or higher, $K\subseteq Y$ is a nontrivial, closed, convex cone with $0\notin\operatorname{cl}\,S_{-K}(||\cdot||)$, $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})$ is a cone in $Y$ with $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$. Moreover, assume that one of the following statements is valid:
  1. a)
    $B_{A}(||\cdot||)$ and ${\mathrm{c}l}\,S_{-K}(||\cdot||)$ are compact.
  2. b)
    $A$ is weakly closed, and either $Y$ is reflexive or ${\mathrm{c}l}\,S_{-K}(||\cdot||)$ and $\operatorname{cl}_{w}\,B_{A}(||\cdot||)$ are weakly compact and $0\notin\operatorname{cl}_{w}\,B_{A}(||\cdot||)$.
Then, there is a norm $||\cdot||^{\prime}\sim||\cdot||$ and a pair $(x^{*},\alpha)\in K^{a\#}(||\cdot||^{\prime})\cap(Y^{*}\times\mathbb{P})$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{||\cdot||^{\prime}}(x^{*},\alpha))$. Moreover, any of the equivalent assertions given in Proposition 5.12, applied for $||\cdot||^{\prime}$ in the role of $\psi$, is valid.
Proof 5.31.
Note that $\varphi_{x^{*},\alpha}$ from the conclusion is constructed with respect to the equivalent norm $||\cdot||^{\prime}$. Since $0\notin\operatorname{cl}\,S_{-K}(||\cdot||)$ and $A\cap\operatorname{cl}(\operatorname{conv}(-K))=A\cap(-K)=\{0\}$, using Proposition 4.11 a) and b) respectively, under our standing assumptions a) and b), one gets that $({\mathrm{c}l}\,S_{A}^{0}(||\cdot||^{\prime}))\cap({\mathrm{c}l}\,S_{-K}(||\cdot||^{\prime}))=\emptyset$ for some norm $||\cdot||^{\prime}\sim||\cdot||$. Applying Lemma 4.9 ($3^{\circ}$, $4^{\circ}$), it follows that ${\mathrm{c}l}\,S_{-K}(||\cdot||^{\prime})$ is compact under a), respectively, weakly compact under b). Consequently, the conclusion follows by applying Theorem 5.25 for $Y$ endowed with the norm $\psi(\cdot):=||\cdot||^{\prime}$ under a), respectively, endowed with the weak topology and $\psi(\cdot):=||\cdot||^{\prime}$ under b).
For the finite dimensional case, we obtain from Corollary 5.30 the following result.
Corollary 5.32.
Assume that ($Y,||\cdot||$) is a real finite dimensional normed space with dimension two or higher, $K\subseteq Y$ is a nontrivial, closed, pointed, convex cone, $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$, and $A:=\mathbb{A}(\bar{x})$ is a closed cone in $Y$ with $\{0\}\neq f[\Omega]-f(\bar{x})\subseteq A$.
Then, the conclusion of Corollary 5.30 holds.
Proof 5.33.
By the well-known Krein-Rutman theorem (see, e.g., [49, Th. 2.4.7]) we know that $K^{\#}\neq\emptyset$, and by [29, Th. 3.1 ($4^{\circ}$)] it follows that $0\notin\operatorname{cl}\,S_{-K}(||\cdot||)$. So, the assumptions of Corollary 5.30 are satisfied.
Remark 5.34.
Besides the proper efficiency concepts in the senses of Benson, Borwein and Hurwicz, we gain also scalarization results for the proper efficiency concept in the sense of Henig taking into account the known relationships between proper efficiency concepts in vector optimization. In order to emphasize this fact, consider a nontrivial cone $K\subseteq Y$ with normlike-base $B_{K}(\psi)$ and suppose that either $K$ is solid or $\psi$ is continuous. Assume that ${\mathrm{P}Eff}_{He}(\Omega\mid f,K)\subseteq{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$. Taking into account Remark 5.27, in this subsection, we have provided sufficient conditions for the equivalence of the following assertions:
  • $1^{\circ}$.
    $\bar{x}\in{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$.
  • $2^{\circ}$.
    $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$.
  • $3^{\circ}$.
    There is $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\operatorname{\mathbb{P}})$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$ (or any of the equivalent assertions given in Proposition 5.12 is valid),
where in a normed setting the third assertion is replaced by
  • $3^{\circ}$’.
    There is a norm $||\cdot||^{\prime}\sim||\cdot||$ and a pair $(x^{*},\alpha)\in K^{a\#}(||\cdot||^{\prime})\cap(Y^{*}\times\mathbb{P})$ such that $\bar{x}\in\operatorname{Eff}(\Omega\mid f,C_{||\cdot||^{\prime}}(x^{*},\alpha))$ (or any of the equivalent assertions given in Proposition 5.12, applied for $||\cdot||^{\prime}$ in the role of $\psi$, is valid).
Remark 5.35.
Under the assumptions in Corollary 5.28 a) (in particular the convexity of $A$), we were able to replace the separation condition (5.2) by $0\notin{\mathrm{c}l}\,S_{-K}$ and $\bar{x}\in{\mathrm{P}Eff}_{\mathbb{A}}(\Omega\mid f,K)$ (which implies the condition $A\cap(-K)=\{0\}$) in our scalarization results. In contrast, the scalarization result (for Henig and Benson proper efficiency concepts) by García-Castaño, Melguizo-Padial and Parzanese [16, Cor. 3.10] in real normed spaces still contains a separation condition (SSP), also under convexity of $A$ (ensured by a convexlikness assumption on $f$).
5.4 Scalarization results for the concept of proper efficiency in the sense of Henig
Let us end the paper with stating some further scalarization results for the concept of proper efficiency in the sense of Henig by making special use of the structure of the dilating cones $D\in\mathcal{D}(K)$.
Lemma 5.36.
Assume that $Y$ is a real locally convex space, $K\subseteq Y$ is a nontrivial cone, and $\bar{D}:=Y\setminus(-\operatorname{int}\,D)$ for some $D\in\mathcal{D}(K)$. Assume that $B_{K}$ and $B_{\bar{D}}$ are normlike-bases for $K$ and $\bar{D}$, respectively, constructed with respect to a continuous seminorm $\psi$. Suppose that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{\bar{D}}^{0}$ is compact, and
$({\mathrm{c}l}\,S_{\bar{D}}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset.$
Then, there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
$K\setminus\{0\}\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)\subseteq\operatorname{int}\,D.$
Proof 5.37.
First, note that $\emptyset\neq K\setminus\{0\}\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)$ for all $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ taking into account the continuity of $\psi$. By Proposition 4.15, there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
$x^{*}(d)+\alpha\psi(d)\geq 0>x^{*}(-k)+\alpha\psi(k)\quad\mbox{for all }d\in\bar{D}\mbox{ and all }k\in K\setminus\{0\}.$ (5.3)
Any $x\in C^{>}_{\psi}(x^{*},\alpha)$ satisfies $0>x^{*}(-x)+\alpha\psi(x)$, hence $-x\notin\bar{D}$ in view of (5.3), and so $x\in\operatorname{int}\,D$.
Theorem 5.38.
Assume that $Y$ is a real locally convex space, and $K\subseteq Y$ is a nontrivial cone. Let $\bar{x}\in{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$, i.e., there is $D\in\mathcal{D}(K)$ such that $\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,D)$. Suppose that $D$ is closed. Define $\bar{D}:=Y\setminus(-\operatorname{int}\,D)$. Assume that $B_{K}$ and $B_{\bar{D}}$ are normlike-bases for $K$ and $\bar{D}$, respectively, constructed with respect to a continuous seminorm $\psi$. Suppose that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{\bar{D}}^{0}$ is compact, and
$({\mathrm{c}l}\,S_{\bar{D}}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset.$
Then, the conclusion of Theorem 5.25 holds.
Proof 5.39.
By Lemma 5.36 there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
$K\setminus\{0\}\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)\subseteq\operatorname{int}\,D.$
Since $D$ is closed, we have $C_{\psi}(x^{*},\alpha)\subseteq D$, hence $C_{\psi}(x^{*},\alpha)\setminus\{0\}\subseteq D\setminus\{0\}$. We conclude that
$\bar{x}\in{\mathrm{E}ff}(\Omega\mid f,D)\subseteq{\mathrm{E}ff}(\Omega\mid f,C_{\psi}(x^{*},\alpha)).$
Finally, noting that $C_{\psi}(x^{*},\alpha)\neq\ell(C_{\psi}(x^{*},\alpha))$ by Lemma 2.11 ($7^{\circ}$), Proposition 5.12 yields the rest of the conclusion.
Theorem 5.40.
Assume that $Y$ is a real locally convex space, and $K\subseteq Y$ is a nontrivial cone. Let $\bar{x}\in{\mathrm{P}Eff}_{He}(\Omega\mid f,K)$, i.e., there is $D\in\mathcal{D}(K)$ such that $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,D)$. Define $\bar{D}:=Y\setminus(-\operatorname{int}\,D)$. Assume that $B_{K}$ and $B_{\bar{D}}$ are normlike-bases for $K$ and $\bar{D}$, respectively, constructed with respect to a continuous seminorm $\psi$. Suppose that one of the sets ${\mathrm{c}l}\,S_{-K}$ and ${\mathrm{c}l}\,S_{\bar{D}}^{0}$ is compact, and
$({\mathrm{c}l}\,S_{\bar{D}}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset.$
Then, there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Moreover, any of the equivalent assertions given in Proposition 5.14 is valid.
Proof 5.41.
By Lemma 5.36 there exists $(x^{*},\alpha)\in K^{a\#}\cap(Y^{*}\times\mathbb{P})$ such that
$K\setminus\{0\}\subseteq C^{>}_{\psi}(x^{*},\alpha)=\operatorname{int}\,C_{\psi}(x^{*},\alpha)\subseteq\operatorname{int}\,D,$
hence $\bar{x}\in{\mathrm{W}Eff}(\Omega\mid f,D)\subseteq{\mathrm{W}Eff}(\Omega\mid f,C_{\psi}(x^{*},\alpha))$. Proposition 5.14 finishes the proof.
Remark 5.42.
Proposition 4.1 gives sufficient conditions for the equality $({\mathrm{c}l}\,S_{\bar{D}}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset$, which is given as an assumption in Theorems 5.38 and 5.40. Suppose that there is $(x^{*},\alpha)\in Y^{*}\times\mathbb{P}$ such that (4.4) is valid. Notice that the set $\bar{D}$ is a nontrivial, closed cone in $Y$ with $\bar{D}\cap(-K)=\{0\}$. Assume that the nontrivial, convex cone $K$ is closed and that either ${\mathrm{c}onv}({\mathrm{c}l}(B_{\bar{D}})\cup\{0\})$ is closed or ${\mathrm{c}l}\,B_{\bar{D}}$ is compact. Then, Proposition 4.1 ensures $({\mathrm{c}l}\,S_{\bar{D}}^{0})\cap({\mathrm{c}l}\,S_{-K})=\emptyset.$
Remark 5.43.
The approach in this Section 5.4 follows basically some ideas by Kasimbeyli [45, Th. 4.4 and 5.7], who deals (in a normed setting) with $\varepsilon$-conic neighborhoods for $K$ (as dilating cones) and supposes a separation property for a whole family of dilating cones.
In Section 5.3 one has a separation condition for $A=\mathbb{A}(\bar{x})$ and $-K$ (for a given $\bar{x}$), and one can further ensure that $A$ is convex. Here in Section 5.4 we have a separation condition for $\bar{D}=Y\setminus(-\operatorname{int}\,D)$ and $-K$. One has to notice that the complement cone $\bar{D}$ is usually not convex if $D$ is convex. Moreover, the cone $\mathbb{A}(\bar{x})$, which has a particular structure for specific choices of $\mathbb{A}$, is more easy to handle than the cone $\bar{D}=Y\setminus(-\operatorname{int}\,D)$ based on a general dilating cone $D\in\mathcal{D}(K)$.
6 Conclusions
It is well-known that several scalarizing functions used in scalarization methods in vector optimization and in the treatment of optimization problems under uncertainty admit certain cone representation and cone monotonicity properties (e.g., methods in the sense of Gerstewitz [20, 21]; Kasimbeyli [45, 46]; Pascoletti and Serafini [53]; Zaffaroni [58]). In the general framework of real topological-linear spaces, our paper provides new insights into the relationships between a vector optimization problem based on an original (convex) cone, an updated vector optimization problem based on an enlarging / dilating cone of Bishop-Phelps type, and scalar optimization problems involving (Bishop-Phelps type) cone representing and cone monotone scalarizing functions. Thus, our new scalarization results with respect to the concepts of weak efficiency and different types of proper efficiency (e.g., in the sense of Benson; Borwein; Henig; Hurwicz) are also useful for several established scalarization methods in vector optimization (see also Remark 5.17).
In future research, we like to derive Bishop-Phelps type scalarization results for approximate solutions of vector optimization problems in real topological-linear spaces, which would extend recent scalarization results obtained by García-Castaño, Melguizo-Padial and Parzanese [16]. Moreover, the derivation of duality statements for such vector optimization problems based on Bishop-Phelps type scalarization is of interest (see also Kasimbeyli and Karimi [47]).